### v1

# KOALA++: Efficient Kalman-Based Optimization of Neural Networks with Gradient-Covariance Products

Abstract
We propose KOALA++, a scalable Kalman-based optimization algorithm that explicitly models structured gradient uncertainty in neural network training. Unlike second-order methods, which rely on expensive second order gradient calculation, our method directly estimates the parameter covariance matrix by recursively updating compact gradient covariance products. This design improves upon the original KOALA framework that assumed diagonal covariance by implicitly capturing richer uncertainty structure without storing the full covariance matrix and avoiding large matrix inversions. Across diverse tasks, including image classification and language modeling, KOALA++ achieves accuracy on par or better than state-of-the-art first- and second-order optimizers while maintaining the efficiency of first-order methods.
1 Introduction
Optimization lies at the heart of deep learning. As neural networks continue to scale in depth, width, and data complexity, the efficiency, stability, and scalability of optimization algorithms have become increasingly important. First-order methods such as stochastic gradient descent (SGD) (robbins1951stochastic, ) and its adaptive variants like Adam (kingma2015adam, ) and RMSProp (tieleman2012lecture, ) offer practical efficiency and have driven much of the progress in deep learning. However, they often suffer from instability, slow convergence, and sensitivity to hyperparameter tuning – particularly in large-scale or noisy training regimes (keskar2017largebatch, ; zhang2019understanding, ).
To mitigate these issues, second-order optimizers such as K-FAC (martens2015optimizing, ) and Shampoo (gupta2018shampoo, ) introduce curvature-aware updates by approximating the Hessian or Fisher information matrix. These methods often improve convergence and generalization but require costly matrix operations, limiting their scalability. Recently, AdaFisher (gomes2024adafisher, ) has emerged as a more scalable alternative, leveraging block-diagonal Kronecker approximations of the Fisher information matrix to reduce overhead while preserving second-order benefits. Despite its effectiveness, AdaFisher fundamentally depends on the quality of the Fisher matrix estimation, which remains sensitive to noise and mini-batch variability.
An alternative and theoretically grounded direction stems from Kalman filtering. In Bayesian estimation theory, the posterior covariance matrix of an unbiased estimator defines the tightest achievable lower bound on parameter uncertainty, as formalized by the Cramér–Rao bound (rao1992information, ; cramer1999mathematical, ). While the Fisher information matrix serves as an inverse bound proxy, it does not directly capture the true uncertainty in parameters. Kalman filters, on the other hand, propagate the full covariance matrix of the estimate through time, providing a more faithful representation of uncertainty – particularly in noisy or nonstationary environments.
Motivated by this insight, KOALA (davtyan2022koala, ) proposed to frame stochastic optimization as a Kalman filtering problem. It models parameter updates as recursive state estimation, using a covariance matrix to adjust step sizes adaptively. However, KOALA’s original formulation relies on overly simplistic approximations, such as diagonal covariances and idealized gradient assumptions, which limit its practical effectiveness.
In this paper, we introduce KOALA++, a principled extension of the KOALA framework that explicitly maintains structured gradient covariance information. Rather than approximating second-order curvature via Fisher matrices, KOALA++ tracks the evolution of uncertainty directly through a recursive update at each training iteration $k$ of the gradient-covariance product $v_{k}=H_{k}P_{k-1}$, where $H_{k}$ is the gradient of the loss function and $P_{k-1}$ is the prior estimate of the parameter covariance. By doing so, KOALA++ effectively avoids expensive matrix inversions and storing the full covariance matrix $P_{k-1}$, while benefiting from the lifted constraints on its structure. Moreover, KOALA++ retains scalability akin to first-order methods, while achieving optimization dynamics and final performance competitive with state-of-the-art second-order methods.
Our main contributions are summarized as follows: We propose KOALA++, a Kalman-inspired optimizer that explicitly propagates structured gradient uncertainty via covariance modeling. We derive an efficient recursive update for the surrogate $v_{k}$ based on a symmetric minimum-norm solution to the covariance estimation problem. We implement KOALA++ with minimal overhead and validate its effectiveness across diverse architectures and tasks, including ResNet family of models he2016deep , vision transformers liu2021swin , and language models. Extensive experiments show that KOALA++ consistently matches or exceeds the accuracy of first- and second-order baselines while offering improved stability and a better efficiency – performance trade-off.
2 Related Work and Motivation
2.1 Optimization Methods: A Brief Overview
First-order optimizers such as SGD robbins1951stochastic and Adam kingma2015adam are widely adopted for their scalability and simplicity, but they often ignore the structure of gradient noise and require manually tuned learning rate schedules.
Second-order optimization methods, including K-FAC martens2015optimizing , Shampoo gupta2018shampoo , and more recently AdaFisher gomes2024adafisher , seek to address these shortcomings by explicitly incorporating curvature information. Nevertheless, practical implementations rely heavily on fixed approximations, such as block-diagonal or Kronecker-factored representations of the Hessian or Fisher matrices, to maintain computational feasibility. Such static approximations fail to adapt dynamically to the non-stationary curvature and noise structure inherent in deep network training.
2.2 Kalman Filtering Perspective for Optimization
To overcome these limitations, recent work proposes modeling optimization as a probabilistic filtering process, inspired by Kalman filtering kalman1960new . In fact, the idea of applying Kalman filtering to neural network optimization is not new. Several earlier works have explored different ways of integrating filtering theory into learning systems. For example, Ismail et al. ismail2018estimation propose a scalar Kalman update rule for stochastic optimization; Shashua and Mannor shashua2019trust apply Kalman filtering in reinforcement learning for value approximation; and Ollivier ollivier2019extended connects the Extended Kalman Filter (EKF) to online natural gradient methods via Amari’s framework. Moreover, de Vilmarest and Wintenberger de2021stochastic apply EKF to logistic regression, and Takedo et al. takenga2004comparison compare GD with Kalman-style variants in control settings.
While these approaches illustrate the versatility of Kalman-based ideas, most are either narrowly scoped to specific domains (e.g., reinforcement learning, linear models), or suffer from limited scalability due to full covariance computation. Among these, KOALA davtyan2022koala stands out as a practical and generalizable solution for deep learning. By interpreting gradient descent as a scalar state estimation problem, KOALA introduces tractable variants that scale to large neural networks using diagonal or isotropic covariance approximations. For a more direct comparison with our method, we report the complete descriptions of KOALA-V and KOALA-M variants in Appendix A.
2.3 Motivation for KOALA++: Towards Structured Covariance Updates
However, the simplifications made in KOALA also constrain their ability to model structured noise or adapt to anisotropic curvature. In particular, KOALA assumes $P_{k}\approx\sigma^{2}I$ and approximates $H_{k}^{\top}H_{k}$ using gradient norms, which may be suboptimal in noisy or high-dimensional settings.
Our proposed extension, KOALA++, addresses these limitations by explicitly tracking a directional covariance projection $v_{k}:=H_{k}P_{k-1}$ and updating it via a recursive formulation of least squares. This approach maintains the efficiency of first-order updates while offering richer uncertainty modeling than prior filtering-based optimizers.
3 KOALA++
3.1 Problem Formulation and Kalman Filtering Notation
The training of neural networks is fundamentally an optimization problem, where given a dataset ${\cal D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, the objective of the learning task is to adjust the model’s parameters $\theta$ such that the network’s predictions $f(x_{i};\theta)$ minimize the empirical risk
$\min_{\theta}{\cal L}(\theta),\quad{\cal L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}l(f(x_{i};\theta),y_{i}),$ (1)
where $l$ is a chosen loss function. However, in modern deep learning applications, $N$ is often large, making full-batch optimization computationally infeasible. As a result, practitioners rely on stochastic optimization methods that operate on minibatches of data, using estimates of the loss
$L_{k}(\theta)=\frac{1}{m}\sum_{i\in B_{k}}l(f(x_{i};\theta),y_{i}),$ (2)
where $B_{k}$ is a set of indices in the $k$-th minibatch and $m=|B_{k}|$ is the minibatch size. While minibatch-based approaches such as SGD robbins1951stochastic and its adaptive variants are effective and scalable, they introduce noise due to the stochastic nature of loss and gradient estimates.
Table 1: EKF equations from davtyan2022koala .
$S_{k}=H_{k}(P_{k-1}+QI)H_{k}^{\top}+R$
$K_{k}=(P_{k-1}+QI)H_{k}^{\top}S_{K}^{-1}$
$P_{k}=(I-K_{k}H_{k})(P_{k-1}+QI)$
$\theta_{k}=\theta_{k-1}+K_{k}(L_{k}^{\text{target}}-L_{k}(\theta_{k-1}))$
To mitigate this issue, a recent approach, KOALA davtyan2022koala , casts neural network training as a recursive state estimation problem using Kalman filtering kalman1960new , a principled Bayesian framework for recursively estimating the hidden state of a system from noisy observations. At each iteration $k$, it combines prior estimates with new measurements to update the state $\theta_{k}$ and its associated uncertainty in the form of a covariance matrix $P_{k}$. Importantly, by keeping track of an estimate of the system’s state’s covariance, Kalman filtering is essentially implicitly using second order information about the state variables. The standard Extended Kalman Filter (EKF) generalizes this to non-linear systems using local linearization. In KOALA, the model parameters $\theta_{k}$ are treated as latent states with stationary evolution under process noise, and the minibatch loss function $L_{k}$ serves as a noisy observation of a target loss value $L_{k}^{\text{target}}$. Under this settings, the EKF equations can simplified as in Table 1. To remain tractable in high-dimensional settings, KOALA simplifies the EKF update by:
Assuming an identity-scaled covariance matrix $P_{k}\approx\sigma_{k}^{2}I$; Approximating $H_{k}^{\top}H_{k}$ with $||H_{k}||^{2}I$; Reducing all matrix operations to scalar computations.
The resulting update closely resembles SGD, but with an adaptively scaled learning rate governed by the uncertainty estimate $P_{k}$.
While KOALA demonstrates promising stability and simplicity, its reliance on scalar approximations limits its capacity to capture directional or structured noise in high-dimensional loss landscapes. To address this, we propose KOALA++, which enhances the original framework with an efficient estimation of directional state covariance while preserving the Kalman-based update structure.
3.2 Kalman Update with $v_{k}$-Based Reparameterization
To overcome the limitations of isotropic covariance assumptions in KOALA, KOALA++ introduces a more expressive formulation that captures directional uncertainty without incurring the full cost of matrix-valued updates. Instead of approximating the full posterior covariance $P_{k}\in\mathbb{R}^{n\times n}$, we propagate a low-rank surrogate
$v_{k}:=H_{k}P_{k-1}\in\mathbb{R}^{1\times n},$ (3)
which represents the projection of the parameter covariance along the current measurement direction. This allows KOALA++ to maintain a richer understanding of uncertainty and curvature structure, while keeping the computational complexity comparable to first-order methods.
Low-Rank Reparameterization of the Kalman Update:
By assuming a constant process noise $Q_{k}=Q$ for all $k$, the innovation covariance $S_{k}$ becomes
$S_{k}$ $=H_{k}(P_{k-1}+Q)H_{k}^{\top}+R=H_{k}v_{k}^{\top}+H_{k}QH_{k}^{\top}+R.$ (4)
Using this expression, the Kalman gain becomes
$K_{k}$ $=\frac{(P_{k-1}+Q)H_{k}^{\top}}{S_{k}}=\frac{v_{k}^{\top}+QH_{k}^{\top}}{S_{k}}$ (5)
and the corresponding state update is
$\theta_{k}\leftarrow\theta_{k}+\left(L_{k}^{\text{target}}-L_{k}(\theta_{k})\right)\frac{v_{k}^{\top}+QH_{k}^{\top}}{S_{k}}.$ (6)
Notice, that as suggested in davtyan2022koala , the learning rate of the update can be controlled by selecting a specific $L_{k}^{\text{target}}$. For instance, $L_{k}^{\text{target}}=(1-\eta_{k})L_{k}(\theta_{k-1})$ ensures that the update is scaled by $\eta_{k}$.
Recursive Update of $v_{k}$:
To avoid storing $P_{k-1}$ explicitly, we recursively update $v_{k}$ by substituting the Kalman covariance update $P_{k-1}=(I-K_{k-1}H_{k-1})(P_{k-2}+Q)$
$v_{k}$ $=H_{k}P_{k-1}=H_{k}(I-K_{k-1}H_{k-1})(P_{k-2}+Q)$
$=(H_{k}-H_{k}\frac{v_{k-1}^{T}+QH_{k-1}^{T}}{S_{k-1}}H_{k-1})(P_{k-2}+Q)$
$=H_{k}P_{k-2}+H_{k}Q-H_{k}\frac{v_{k-1}^{T}+QH_{k-1}^{T}}{S_{k-1}}v_{k-1}-H_{k}\frac{v_{k-1}^{T}+QH_{k-1}^{T}}{S_{k-1}}H_{k-1}Q.$ (7)
For simplicity, we define
$\lambda_{k}$ $=H_{k}\frac{v_{k-1}^{T}+QH_{k-1}^{T}}{S_{k-1}},$ (8)
which allows to rewrite the update for $v_{k}$ in a more compact way:
$v_{k}=H_{k}P_{k-2}-\lambda_{k}v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q.$ (9)
In the last equation, all the terms except for the first one can be efficiently calculated at iteration $k$. The first term instead cannot be easily evaluated due to the $P_{k-2}$ term. The objective is to never store it explicitly to avoid the quadratic storage demand. In principle, one can approximate $H_{k}$ with $H_{k-1}$ and replace the problematic term with $v_{k-1}$. However, we observed experimentally that this does not work well due to the noise in the gradients.
To proceed, we aim to retrospectively approximate the prior covariance matrix $P_{k-2}$ instead. We do this by leveraging the relation $v_{k-1}=H_{k-1}P_{k-2}$. However, since this equation is generally underdetermined in high-dimensional settings, we formulate a regularized inverse problem to obtain a unique and tractable solution. Specifically, we adopt the minimum Frobenius-norm formulation
$\min_{P_{k-2}}\|P_{k-2}\|_{F}^{2}\quad\text{subject to}\quad H_{k-1}P_{k-2}=v_{k-1},$ (10)
which yields a closed-form least-squares solution. This construction allows us to approximate $P_{k-2}$ efficiently without explicitly storing or inverting full covariance matrices.
For computational and memory efficiency, we use only one linear constraint, $v_{k-1}=H_{k-1}P_{k-2}$, although more could be used in principle. However, introducing more constraints, e.g. projections of other past gradients $H_{k-2},H_{k-3}$ etc., would also require to memorize more gradient terms and using more calculations to update $v_{k}$. Finally, the use of the minimum Frobenius-norm criterion is to ensure that $P_{k-2}$ has a bounded norm and thus a more stable update for $v_{k}$.
In the next subsection we discuss two variants of our algorithm, depending on the nature of the solution to the optimization problem (10).
3.3 Least-Squares Estimate of $P_{k-2}$
Variant I: Vanilla Least-Squares Estimation
In the first variant of our approach, we solve the least-squares problem without imposing any structural constraints on the covariance estimate. Given the relation $v_{k-1}=H_{k-1}P_{k-2}$, we compute $P_{k-2}\in\mathbb{R}^{n\times n}$ by introducing Lagrange multipliers $\lambda\in\mathbb{R}^{1\times n}$ and write the Lagrangian
$\mathcal{L}(P_{k-2},\lambda)=\|P_{k-2}\|_{F}^{2}+(H_{k-1}P_{k-2}-v_{k-1})\lambda^{T}.$ (11)
This formulation admits a closed-form solution [^1] See Appendix B.1 for a detailed derivation., which is:
$P_{k-2}=\frac{H_{k-1}^{T}v_{k-1}}{||H_{k-1}||^{2}}.$ (12)
By substituting this solution into $H_{k}P_{k-2}$ gives
$H_{k}P_{k-2}=\alpha_{k}v_{k-1},\quad\text{where}\quad\alpha_{k}:=\frac{H_{k}H_{k-1}^{T}}{\|H_{k-1}\|^{2}}.$ (13)
Finally, we can get the vanilla $v_{k}$ update as follows
$v_{k}=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q.$ (14)
Variant II: Symmetric Covariance Estimation
To improve stability and theoretical consistency with classical Kalman filtering, we introduce a second variant that enforces symmetry in the estimation of $P_{k-2}$. In particular, the additional regularization term $P_{k-2}=P_{k-2}^{T}$ is added to the objective. As in the previous derivation, we obtain a closed-form solution for $P_{k-2}$ [^2] See Appendix B.2 for a detailed derivation.
$P_{k-2}=\frac{H_{k-1}^{\top}\,v_{k-1}+v_{k-1}^{\top}\,H_{k-1}}{\|H_{k-1}\|^{2}}-H_{k-1}\,v_{k-1}^{\top}\,\frac{H_{k-1}^{\top}\,H_{k-1}}{\|H_{k-1}\|^{4}}.$ (15)
When substituting this into $H_{k}P_{k-2}$, the only difference between the two variants is the term
$w_{k}$ $=\frac{(H_{k}v_{k-1}^{T})(H_{k-1}H_{k-1}^{T})-(H_{k-1}v_{k-1}^{T})(H_{k}H_{k-1}^{T})}{\|H_{k-1}\|^{4}}.$ (16)
To simplify notation and implementation, we adopt a unified update rule for both variants. In the unconstrained case, we simply set $w_{k}=0$, effectively reducing the update to the asymmetric version. This allows the general form of the update to be written compactly as
$\boxed{v_{k}=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q+w_{k}H_{k-1}.}$ (17)
To maintain consistency with the original KOALA formulation and to simplify computation in the early stages of training, we initialize the directional covariance product $v_{1}=H_{1}P_{0}$ using a scaled identity assumption on the initial covariance matrix $P_{0}$. Specifically, we set:
$v_{1}:=\sigma_{0}^{2}H_{1},$ (18)
where $\sigma_{0}^{2}$ is a scalar hyperparameter representing the initial uncertainty. This corresponds to assuming $P_{0}=\sigma_{0}^{2}I$, which aligns with the isotropic prior used in KOALA and allows us to avoid explicit matrix computations at initialization. The general procedure of KOALA++ is summarized in Algorithm 1.
Algorithm 1 KOALA++
  Initialize $\theta_{0}$, $v_{1}$, $Q$, $R$, and fix the learning rate schedule $\eta_{k}$
  for $k=2$ to $T$ do
     For simplicity, denote $H_{k}=\nabla L_{k}(\theta_{k-1})$Calculate $\alpha_{k},\lambda_{k},w_{k}$ respectively from Equations (13) (8) (16)
     Update:
$v_{k}$ $=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q+w_{k}H_{k-1}$ (19)
$\theta_{k}$ $=\theta_{k-1}-\frac{\eta_{k}L_{k}(\theta_{k-1})}{H_{k}v_{k}^{T}+H_{k}QH_{k}^{T}+R}\cdot(v_{k}^{T}+QH_{k}^{T})$ (20)
  end for
  return  $\theta_{T}$
4 Experiments
We evaluate KOALA++ on a range of vision and language tasks to demonstrate its generality, stability, and convergence behavior compared to existing optimizers. Our experiments are organized into three parts: image classification, language modeling and ablation studies. Due to space constraints, we report a representative subset of results in the main text. Additional evaluations, including ablations, hyperparameter tuning, and extended visualizations, are provided in Appendix D.
4.1 Image Classification
CIFAR-10/100 Classification:
We follow the experimental setup of the original KOALA paper davtyan2022koala for CIFAR-10 and CIFAR-100 classification tasks, including data augmentation, model architectures, and optimization settings. We introduce two modifications: (i) training is repeated with three random seeds (42, 3407, and 2025) to ensure statistical robustness, and (ii) we adopt a two-stage learning rate decay strategy at the 100th and 150th epochs for 200-epoch training (inspired by successful multi-step decay strategies reported in prior works such as rice2020overfitting ), which improves generalization compared to the single-step decay used in the original paper. The results are shown in Table  2.
Table 2: Experimental results on CIFAR-10 and CIFAR-100 with different optimizers for 100 and 200 epochs. Best and second best performance is indicated in boldface and with a blue shading.
Dataset Architecture Method Error
100-epochs 200-epochs
Top-1 Err Top-5 Err Top-1 Err Top-5 Err
[…]
Adam 7.09 0.30 6.09 0.29
KOALA-M 6.34 0.30 5.85 0.25
[…]
Adam 6.43 0.27 5.74 0.33
KOALA-M 8.00 0.24 8.16 0.33
[…]
KOALA-M 7.62 0.30 7.67 0.32
[…]
CIFAR-100 ResNet-50 SGD 25.05 6.74 22.06 5.71
Adam 24.95 6.96 24.44 6.81
[…]
W-ResNet-50-2 SGD 23.83 6.35 22.47 5.96
Adam 23.73 6.64 24.07 7.06
[…]
Comparison with Recent Advances
To comprehensively evaluate the effectiveness and generalization capability of KOALA++, we design two experimental settings that progressively incorporate more recent optimizers, modern network architectures, stronger learning rate schedulers, and richer data augmentations. This design mirrors recent benchmarking practices in the literature, particularly those from KOALA davtyan2022koala and AdaFisher gomes2024adafisher .
(1) CIFAR-100 + ResNet-50 + Advanced Optimizers. To evaluate KOALA++ in a standardized yet competitive setting, we follow the benchmarking protocol introduced in KOALA davtyan2022koala , and compare our method with a wide range of recent optimizers on the CIFAR-100 dataset using ResNet-50, trained for 100 epochs. All optimizers are tested under the same learning rate schedule (StepLR) and training configuration. For each optimizer, we adopt hyperparameters as suggested in the original papers or official implementations. The settings and results are shown in Table 3
Table 3: CIFAR-100 (ResNet-50, 100 epochs, StepLR) results and optimizer hyperparameters.
Optimizer Top-1 Err. Top-5 Err. Hyperparameters / Source
Yogi reddi2018adaptive 33.99 10.90 lr=$10^{-2}$, $\beta_{1}$=0.9, $\beta_{2}$=0.999, $\epsilon$=$10^{-3}$
Adamax kingma2015adam 32.42 10.74 Same as Adam
AdamW loshchilov2017decoupled 27.23 7.98 Same as Adam
AdamP heo2020adamp 26.62 7.61 Same as Adam
Amsgrad reddi2019convergence 25.27 6.78 Same as Adam
Fromage bernstein2020distance 24.65 6.71 lr=$10^{-2}$ (recommended in GitHub [^3] https://github.com/jxbz/fromage#voulez-vous-du-fromage)
Adabelief zhuang2020adabelief 23.07 6.05 Official config [^4] https://github.com/juntang-zhuang/Adabelief-Optimizer#
Adan xie2024adan 24.92 6.86 Paper and repo defaults [^5] https://github.com/lucidrains/Adan-pytorch
AdaFisher gomes2024adafisher 23.38 6.05 Original AdaFisher hyperparams
[…]
(a) Test Loss (log)
(b) Test Top-1 Error
Figure 1: Comparison of training loss, validation loss, and validation error for different optimizers on CIFAR-100 using ResNet-50. KOALA++ demonstrates the strongest performance drop at scheduled learning rate decays (epochs 30, 60, and 90), highlighting its superior scheduler responsiveness.
Figure 1 presents the training and validation dynamics on CIFAR-100 with ResNet-50 under a multi-step learning rate schedule at epochs 30, 60, and 90. Compared to SGD, Adam, AdaFisher, and KOALA-M, our proposed KOALA++ demonstrates consistently lower validation loss in log-scale, along with improved generalization as shown by the lower Top-1 error.
In particular, the benefits of KOALA++ are most evident as the scheduled learning rate drops: the optimizer exhibits sharp improvements in both loss and error immediately following the milestones, indicating more effective adaptation to the changing optimization dynamics. This suggests that KOALA++ better integrates structural information from past gradients and uncertainty estimates, allowing it to respond more effectively to scheduler-induced shifts.
Moreover, KOALA++ maintains competitive or superior stability across all training phases, without exhibiting the overfitting spikes or stagnation observed in other optimizers. These observations collectively validate KOALA++ as a robust and efficient optimizer for large-scale vision tasks.
(2) CIFAR-10 + Modern Models + Cosine Annealing + Strong Augmentation To align with the stronger benchmark protocol proposed by AdaFisher gomes2024adafisher , we conduct a second set of experiments that adopt more modern practices in deep learning optimization:
  • Advanced models: ResNet-101, DenseNet-121 huang2017densely , MobileNetV3 howard2019searching , and the Swin Transformer(Tiny) liu2021swin , reflecting a diversity of architectures including convolutional and attention-based designs.
  • Modern learning rate scheduler: We use CosineAnnealingLR for all optimizers, which has been shown to improve convergence stability and final accuracy in modern setups.
  • Stronger data augmentation: In addition to standard cropping and flipping, we apply Cutout devries2017improved during training to improve generalization.
  • Larger batch size: Following current best practices, we increase the batch size to 256, which can benefit optimizers that are sensitive to noisy gradients.
  • More Random Seeds: Following the AdaFisher benchmark protocol gomes2024adafisher , we report top-1 accuracy averaged over five random seeds: 42, 3407, 9331, 7, and 2025. The standard deviation across these runs is shown in the bottom-right corner of each table entry.
Table 4: Top-1 accuracy (%) on CIFAR-10 across various architectures. We report the average over 5 seeds. The large number is the mean and the subscript is the standard deviation.
Optimizer ResNet50 ResNet101 DenseNet121 MobileNetV3 Tiny Swin
SGD 95.710.1 95.980.2 96.090.1 94.430.2 82.340.2
Adam 94.450.2 94.570.1 94.860.1 93.320.1 87.370.6
AdaHessian 95.540.1 95.290.6 96.110.1 92.860.1 84.150.2
K-FAC 95.660.1 96.010.1 96.120.1 94.340.1 64.790.5
Shampoo 94.590.1 94.630.1 95.660.1 93.810.2 63.910.4
[…]
As shown in Table 4, KOALA++ performs on par with AdaFisher across a range of architectures, while exhibiting consistently lower standard deviations, indicating greater stability across random seeds. All KOALA++ results were obtained without extensive hyperparameter adjustment, highlighting the robustness of the method in the default settings. It is also worth noting that on Swin-Tiny, KOALA++ reaches strong performance with significantly smaller variance, demonstrating better adaptability to Transformer-based architectures.
4.2 Language Model
We evaluate KOALA++ on the Wikitext-2 dataset, following the same language modeling benchmark setting introduced in the AdaFisher paper gomes2024adafisher . The model is a scaled-down version of GPT-1 with four masked self-attention layers and approximately 28 million parameters. We adopt the same architecture, tokenizer, and 50-epoch training schedule as used in the original benchmark.
In line with the original AdaFisher benchmark (See Github [^6] https://github.com/AtlasAnalyticsLab/AdaFisher/tree/main), we report the test perplexity (PPL) corresponding to the best-performing model on the validation set, along with the total training time, to provide a holistic view of performance and efficiency. The results are summarized in Table 5. In addition to reporting test perplexity and total training time, we further provide visualizations to better illustrate the training dynamics(See Appendix D)
4.3 Ablation Study
Due to space constraints, we report only the core ablation comparing KOALA-M and two variants of KOALA++ on CIFAR-100 with ResNet-50. The default KOALA++ uses a symmetric gain approximation, while KOALA++ (NS) adopts an asymmetric covariance. As shown in Figures 2, KOALA++ achieves faster convergence and better final performance. KOALA++ (NS) is competitive but less stable in later stages, underscoring the benefit of structural symmetry. Full ablations are provided in Appendix C.
(a) Test Loss (log)
(b) Test Top-1 Error
Figure 2: Ablation study comparing KOALA++, its asymmetric variant KOALA++ (NS), and KOALA-M on CIFAR-100 with ResNet-50. Left: Log-scale test loss curves. Right: Validation Top-1 error rates.
Table 5: Language modeling performance on Wikitext-2. We report the test perplexity (lower is better) corresponding to the best validation model, as well as the total training time over 50 epochs.
Optimizer Test PPL ↓ Average Training Time per Epoch (s)
AdamW 173.98 23.11
AdaHessian 407.69 62.91
[…]
5 Conclusions, Limitations and Future Work
We proposed KOALA++, a principled and efficient optimization framework that extends the Kalman-based KOALA algorithm by incorporating a lightweight directional covariance approximation. KOALA++ retains the stability and simplicity of its predecessor while substantially improving its capacity to capture structured uncertainty in high-dimensional loss landscapes. Through extensive experiments across vision and language tasks, we demonstrate that KOALA++ achieves competitive or superior performance compared to both first- and second-order optimizers, while maintaining training efficiency on par with standard first-order methods.
Limitations. A notable limitation of our method is that the approximated covariance matrix $P_{k}$ is not explicitly constrained to be positive semi-definite (PSD). Although we empirically observe that the eigenvalues of $P_{k}$ remain stable and bounded throughout training (see Appendix B.3), the lack of a formal PSD guarantee may pose risks in certain settings or applications involving sensitive numerical stability. Addressing this limitation efficiently is left to future work.
Future Work. To address this issue, future work may explore enforcing PSD constraints via projection, spectral regularization, or reparameterization techniques. Additionally, we plan to investigate the integration of KOALA++ with adaptive gradient clipping and apply it to large-scale pretraining scenarios such as vision-language models, diffusion models or instruction-tuned LLMs.
Acknowledgements
Calculations were performed on UBELIX (https://www.id.unibe.ch/hpc), the HPC cluster at the University of Bern. Aram Davtyan has been supported by SNSF Grant 10001278.
Appendix Contents
[…]
1
Appendix A KOALA
In this section, we are going to revisit the original KOALA framework, which was only briefly introduced in the main text. KOALA [3] is a Kalman-inspired optimization algorithm that interprets gradient-based updates as a recursive state estimation process. To ensure scalability in high-dimensional deep learning settings, the original work proposed two tractable variants: KOALA-V and KOALA-M. These variants differ in how they approximate the parameter covariance and in whether they incorporate momentum dynamics. In what follows, we provide a self-contained and detailed exposition of both KOALA-V and KOALA-M, including their derivations, update rules, and practical considerations. This not only offers a clearer understanding of KOALA as a foundational method, but also sets the stage for the improvements introduced in KOALA++.
A.1 Risk Minimization as Loss Adaptivity
The core idea behind KOALA is to cast the optimization of deep neural networks as a state estimation problem under uncertainty. Specifically, the training of neural networks is considered as minimizing a noisy risk function ${L}_{k}$, which is a stochastic estimate of the empirical risk ${\cal L}$. Motivated by the central limit theorem, KOALA assumes that the mini-batch risk ${L}_{k}$ can be modeled as a Gaussian random variable with scalar noise:
${L}_{k}(\theta)\simeq{\cal L}(\theta)-v_{k},\quad\text{where }v_{k}\sim\mathcal{N}(0,R_{k}),$ (21)
where $\theta$ denotes the model parameters and $R_{k}$ is the variance of the observation noise.
Rather than optimizing $\min_{\theta}{\cal L}(\theta)$ directly, KOALA reformulates the problem as adapting the model parameters such that the stochastic risk ${L}_{k}(\theta_{k})$ matches a desired target risk ${L}^{\text{target}}_{k}$, i.e.,
${L}_{k}(\theta_{k})={L}^{\text{target}}_{k}-v_{k}.$ (22)
This formulation enables the use of Kalman filtering, which provides a principled way to recursively update the estimate of $\theta_{k}$ by fusing prior knowledge with new, noisy observations. The underlying system can be described as a standard dynamical model:
$\theta_{k}$ $=\theta_{k-1}+w_{k-1},\quad w_{k}\sim\mathcal{N}(0,Q_{k})$ (23)
${L}^{\text{target}}_{k}$ $={L}_{k}(\theta_{k})+v_{k},\quad v_{k}\sim\mathcal{N}(0,R_{k})$ (24)
A.2 KOALA-V (Vanilla)
The idea behind KOALA-V is remarkably natural, which is directly appling the standard Kalman filtering update equations to the system in Equations (23) and (24). However, directly computing the posterior covariance $P_{k}\in\mathbb{R}^{n\times n}$ is infeasible for high-dimensional models. To address this, KOALA-V introduces the following approximations:
  • The posterior covariance $P_{k}$ is approximated by a scaled identity matrix:
    $P_{k}\approx\sigma_{k}^{2}I_{n}.$ (25)
  • $H_{k}^{\top}H_{k}$ is approximated by the squared norm of the gradient:
    $H_{k}^{\top}H_{k}\approx\left\|\nabla_{\theta}{L}_{k}(\theta_{k})\right\|^{2}I_{n}.$ (26)
The resulting parameter update rule is given in Algorithm 2.
This update closely resembles the SGD update:
$\theta_{k}=\theta_{k-1}-\eta\nabla\hat{L}_{k}(\theta_{k}),$ (27)
but with a key difference: in KOALA-V, the scaling of the gradient is automatically adapted based on the current loss, gradient norm, and covariance estimate. This adaptivity makes KOALA-V more robust to noisy training dynamics than static first-order methods like SGD or Adam.
Algorithm 2 KOALA-V (Vanilla)
  Initialize $\theta_{0}$, $P_{0}$, $Q$ and $R$
  for $k$ in range(1, $T$do
     Predict:
$\hat{\theta}_{k}=\theta_{k-1};\quad\hat{P}_{k}=P_{k-1}+Q$
     Update:
$\theta_{k}$ $=\hat{\theta}_{k}-\frac{\hat{P}_{k}(\hat{L}_{k}(\hat{\theta}_{k})-\hat{L}_{k}^{\text{target}})}{\hat{P}_{k}\|\nabla\hat{L}_{k}(\hat{\theta}_{k})\|^{2}+R}\cdot\nabla\hat{L}_{k}(\hat{\theta}_{k})$ (28)
$P_{k}$ $=\frac{R}{\hat{P}_{k}\|\nabla\hat{L}_{k}(\hat{\theta}_{k})\|^{2}+R}\cdot\hat{P}_{k}$ (29)
  end for
  return $\theta_{T}$
A.3 KOALA-M (Incorporating Momentum Dynamics)
KOALA-M extends the original KOALA framework by explicitly integrating momentum-based dynamics into the Kalman filtering formulation. Instead of modeling the parameter update as a simple random walk, KOALA-M introduces an auxiliary momentum state $p_{k}$, resulting in the augmented state dynamics:
$\theta_{k}$ $=\theta_{k-1}+p_{k-1}+w_{k-1},\quad w_{k-1}\sim\mathcal{N}(0,Q),$ (30)
$p_{k}$ $=\kappa\,p_{k-1}+u_{k-1},\quad u_{k-1}\sim\mathcal{N}(0,U),$ (31)
where $\kappa$ controls the momentum persistence. This momentum augmentation enables KOALA-M to better capture and leverage correlations across sequential updates, effectively enhancing stability.
However, KOALA-M inherits computational complexity due to its augmented state. To mitigate this, KOALA-M approximates the posterior covariance with a structured form, typically assuming it to be a Kronecker product of a $2\times 2$ covariance matrix with an identity matrix $I_{n}$. Similar to KOALA-V, KOALA-M employs layer-wise covariance updates to obtain a better approximation of parameter uncertainty, partially alleviating the limitation of overly simplistic covariance structures.
In contrast, our proposed KOALA++ algorithm does not require explicit momentum augmentation or restrictive covariance approximations due to its inherently richer and adaptive covariance representation. KOALA++ directly maintains structured gradient-covariance products without imposing additional structural assumptions on the covariance matrix $P$, thus naturally capturing both directional information and momentum-like dynamics implicitly.
Appendix B Mathematical Foundations of KOALA++
In the main paper, we provided both the non-symmetric and symmetric solution for $P_{k-2}$. In this section, we provide the detailed derivations of those. For notational simplicity, where possible we drop the subscript in $k$, and we can restate the optimization problem as follows:
$\min_{P}\|P\|_{F}^{2}\quad\text{subject to}\quad HP=v$ (32)
B.1 Vanilla Optimal Solution of Covariance Matrix P
To solve this constrained optimization problem, we introduce a Lagrange multiplier $\lambda\in\mathbb{R}^{1\times n}$ and construct the corresponding Lagrangian function:
$\mathcal{L}(P,\lambda)=\|P\|_{F}^{2}+\lambda(HP-v)^{\top}.$ (33)
We then find the optimal solution by differentiating the Lagrangian $\mathcal{L}$ with respect to $P$, and set this derivative to zero to find critical points:
$\frac{\partial\mathcal{L}(P,\lambda)}{\partial P}=2P+H^{\top}\lambda=0\quad\Rightarrow\quad P=-\frac{1}{2}H^{\top}\lambda.$ (34)
Now, substituting this optimality condition into the constraint equation $HP=v$, we have:
$H\left(-\frac{1}{2}H^{\top}\lambda\right)$ $=v\Rightarrow\lambda=-2\frac{v}{HH^{\top}}.$ (35)
Thus, the optimal solution for $P$ is obtained by substituting back the expression for $\lambda$:
$\boxed{P=\frac{H^{\top}v}{HH^{\top}}}.$ (36)
This solution is referred to as the vanilla optimal solution (asymmetric) because it does not impose symmetry constraints explicitly on $P$. In the next subsection, we derive a symmetric solution variant that respects the symmetry requirement typically imposed on covariance matrices.
B.2 Symmetric Optimal Solution of Covariance Matrix P
For derivation of the symmetric solution, we use the same Lagrangian framework introduced earlier, but now explicitly incorporate the symmetry constraint into the formulation. The Lagrangian function, considering the symmetry constraint, remains the same as before (see Equation (33)).
We use the symmetric matrix derivative form from the Matrix Cookbook [20].Taking the derivative with respect to $P$ and setting it to zero yields
$\frac{\partial\mathcal{L}}{\partial P}(\text{Symmetric})=\left[\frac{\partial\mathcal{L}}{\partial P}\right]+\left[\frac{\partial\mathcal{L}}{\partial P}\right]^{T}-\text{diag}\left[\frac{\partial\mathcal{L}}{\partial P}\right]=0$ (37)
By plugging in Equation (33), we have
$02P+\lambda^{T}H+2P^{T}+H^{T}\lambda-\text{diag}(2P+\lambda^{T}H)=$ 0 (38)
Since $P$ is symmetric, we can rewrite the Equation (38) as:
$4P+\lambda^{T}H+H^{T}\lambda-\text{diag}(2P+\lambda^{T}H)=0,$ (39)
which gives the equations as follows:
$\begin{cases}4P+\lambda^{\top}H+H^{\top}\lambda=0&\text{(off diagonal)}\\ 2P+H^{\top}\lambda=0&\text{(on diagonal)}\end{cases}$ (40)
We observe that the off-diagonal solution satisfies the on-diagonal equation on the diagonal. Therefore, we only need to solve the off-diagonal case. From the off-diagonal constraint $04P+\lambda^{\top}H+H^{\top}\lambda=$ 0, we have $P=-\frac{1}{4}\left(\lambda^{\top}H+H^{\top}\lambda\right)$. By substituting in $HP=v$, we obtain
$HP=-\frac{1}{4}(H\lambda^{\top}H+HH^{\top}\lambda)=v.$ (41)
We can solve $\lambda$ as
$\lambda=-4v(H^{\top}H+HH^{\top}I)^{-1},$ (42)
where $I$ denotes the identity matrix. To compute the inverse, we recall the Woodbury matrix identity [20]:
$(A+UV)^{-1}=A^{-1}-A^{-1}U(I+VA^{-1}U)^{-1}VA^{-1}$ (43)
By letting $A=HH^{\top}I$, $U=H^{\top}$, and $V=H$, we find that:
$(HH^{\top}I+H^{\top}H)^{-1}=\frac{1}{HH^{\top}}\left(I-H^{\top}\cdot\frac{1}{2}H\cdot\frac{1}{HH^{\top}}\right)$ (44)
This provides an expression for $\lambda$ without matrix inverses, i.e.,
$\lambda=-4v\frac{1}{HH^{\top}}\left(I-\frac{1}{2}\frac{H^{\top}H}{HH^{\top}}\right).$ (45)
Finally, we can plug this expression to have a closed-form optimal solution $P^{*}$, that is,
$P^{*}$ $=\frac{1}{HH^{\top}}\left(I-\frac{1}{2}\frac{H^{\top}H}{HH^{\top}}\right)v^{\top}H+H^{\top}v\frac{1}{HH^{\top}}\left(I-\frac{1}{2}\frac{H^{\top}H}{HH^{\top}}\right)$
$=\frac{v^{\top}H+H^{\top}v}{HH^{\top}}-\frac{Hv^{\top}}{(HH^{\top})^{2}}H^{\top}H.$ (46)
B.3 Eigenvalues Analysis of the Optimal Matrix $P^{*}$
Recall the closed-form expression for the optimal matrix
$P^{*}=\frac{1}{HH^{T}}(H^{\top}v+v^{T}H)-\frac{Hv^{T}}{(HH^{T})^{2}}H^{T}H.$ (47)
To analyze the definiteness of $P^{*}$, we consider its eigenvalues. Let $W\in\mathbb{R}^{n\times 1}$ be an eigenvector of $P^{*}$, satisfying
$\lambda W=P^{*}W$ (48)
We can immediately see that $W$ lies in the span of $H^{\top}$ and $v^{\top}$, because the range of $P$ lies in their span. Then, we can write
$W=\alpha H^{\top}+\beta v^{\top}.$ (49)
By substituting it into the eigenvalue equation we obtain
$\lambda(\alpha H^{\top}+\beta v^{\top})=\frac{1}{HH^{\top}}(H^{\top}v+v^{\top}H)(\alpha H^{\top}+\beta v^{\top})-\frac{Hv^{\top}}{(HH^{\top})^{2}}H^{\top}H(\alpha H^{\top}+\beta v^{\top}).$ (50)
We now compute both terms. For simplicity, let us name the following three scalar values as
$A=HH^{\top},\quad B=Hv^{\top},\quad C=vv^{\top}.$ (51)
Then, we have
$P^{*}W$ $=\frac{1}{A}(H^{T}v\alpha H^{T}+H^{T}v\beta v^{T}+v^{T}H\alpha H^{T}+v^{T}H\beta v^{T})$ (52)
$-\frac{1}{A^{2}}(Hv^{T}H^{T}H\alpha H^{T}+Hv^{T}H^{T}H\beta v^{T})$ (53)
$=\frac{1}{A}(\alpha BH^{T}+\beta CH^{T}+\alpha Av^{T}+\beta Bv^{T})-\frac{1}{A^{2}}\alpha BAH^{T}-\frac{1}{A^{2}}\beta B^{2}H^{T}$ (54)
$=\frac{\beta C}{A}H^{T}+\alpha v^{T}+\frac{\beta B}{A}v^{T}-\frac{\beta B^{2}}{A^{2}}H^{T}$ (55)
$=\frac{\beta(AC-B^{2})}{A^{2}}H^{T}+\frac{\alpha A+\beta B}{A}v^{T}.$ (56)
By matching both sides coefficient-wise, we derive a 2D eigenvalue system as follows
$\alpha\lambda$ $=\dfrac{\beta(AC-B^{2})}{A^{2}}$ (57)
$\beta\lambda$ $=\dfrac{\alpha A+\beta B}{A}.$ (58)
If $\alpha=0$, then from Equation (57) we know that $\beta=0$ or $AC-B^{2}=0$. However, $\beta=0$ implies that $W=\alpha H^{\top}+\beta v^{\top}=0$, which contradicts the assumption that $W$ is a non-zero eigenvector. On the other hand, by the Cauchy-Schwarz inequality, $AC-B^{2}=0$ implies that the two directions $H^{\top}$ and $v^{\top}$ become linearly dependent. In this case, the eigenvalues of $P^{*}$ are 00 and $a$, where $H^{\top}=av^{\top}$, due to the linear dependence. The linear dependence is, however, very unlikely in practice, as, in general $H$ and $v$ span a two-dimensional subspace.
Therefore, we must have $\alpha\neq 0$. In this case, the eigenvalue $\lambda$ must satisfy
$\lambda=\frac{\beta}{\alpha}\cdot\frac{AC-B^{2}}{A^{2}}.$ (59)
Now, we need to solve the linear system to examine the sign of $\frac{\beta}{\alpha}$
$(AC-B^{2})\left(\frac{\beta}{\alpha}\right)^{2}-AB\frac{\beta}{\alpha}-A^{2}=0.$ (60)
We get
$\frac{\beta}{\alpha}=\frac{AB\pm\sqrt{A^{2}B^{2}+4(AC-B^{2})A^{2}}}{2(AC-B^{2})}.$ (61)
Finally, the eigenvalues of the matrix $P$ are
$\lambda_{1,2}=\frac{AB\pm\sqrt{A^{2}B^{2}+4(AC-B^{2})A^{2}}}{2A^{2}}.$ (62)
Although the matrix $P$ obtained through least-squares minimization does not guarantee positive definiteness, we conducted an empirical analysis by visualizing the evolution of its eigenvalues at different training stages.
Specifically, we performed eigenvalue tracking on the ResNet-50 model trained on CIFAR-100 at 100 epochs. The learning rate was decayed at epochs 30, 60, and 90, which correspond to the early, middle, and late stages of training, respectively.
To better understand the behavior of the matrix $P$, we analyzed the distribution of its eigenvalues across all mini-batches at four critical epochs: 0, 30, 60, and 90. The results are shown in Figure 3. We found the cases for 60 and 90 epochs to give identical trends as for 30 epochs case. Thus, to avoid unnecessary repetition, we are not showing the plots for these cases.
At epoch 0 (Figure 3, left), the eigenvalues exhibit a wide dynamic range with both large and small eigenvalues increasing significantly over time. The large eigenvalue (orange) and small eigenvalue (blue) rise rapidly and display sharp oscillations, indicating an evolving optimization landscape in the early stage of training.
However, by epoch 30 (Figure 3, right), the eigenvalue trajectories stabilize, and the magnitude of fluctuations reduces dramatically. From epoch 30 onwards, the eigenvalue distributions become increasingly similar. In these later epochs, both the large and small eigenvalues remain relatively flat and nearly constant throughout training, reflecting a stabilized matrix structure and learning dynamics.
(a) Epoch 0
(b) Epoch 30
Figure 3: Evolution of the large and small eigenvalues of the matrix $P$ at epoch 0 and epoch 30 across all mini-batches.
Appendix C Ablations
C.1 Effect of Batch Size on KOALA++ Performance
We evaluate the impact of batch size on the performance of KOALA++ using ResNet50 on the CIFAR100 dataset. As shown in Table 6, increasing the batch size leads to a noticeable degradation in both Top-1 and Top-5 accuracy. The results suggest that KOALA++ is more effective in smaller-batch training regimes, possibly due to better gradient variance estimation and more frequent updates.
Batch Size Top-1 Error (%) Top-5 Error (%)
64 20.27 4.91
128 20.38 4.76
256 21.78 5.84
512 23.63 6.59
Table 6: Impact of batch size on KOALA++ performance (ResNet50, CIFAR100).
C.2 Different Schedulers
In this experiment, we evaluate the impact of different learning rate schedulers on the performance of KOALA++ using the CIFAR-100 dataset. We consider two representative architectures: ResNet-50 as a typical CNN-based model and Swin-Tiny as a typical ViT-based model. All models are trained for 200 epochs using the same optimizer (KOALA++) and hyperparameters.
The partially recorded results are summarized in Table 7.
Table 7: Top-1 and Top-5 error (%) on CIFAR-100 using KOALA++ with different schedulers.
Schedulers ResNet-50 (CNN) Swin-Tiny (ViT)
Top-1 Error Top-5 Error Top-1 Error Top-5 Error
Multi-step 21.80 5.89
CosineAnnealingLR 20.57 5.07 35.05 12.07
Warmup + Cosine 32.18 10.64
As Table 7 shows, different schedulers can lead to notable differences in final top-1 and top-5 error. We now take a closer look at how these schedulers influence the test loss dynamics during training.
For ResNet-50, we compare the Multi-step scheduler and the CosineAnnealingLR scheduler. As shown in Figure 4, CosineAnnealingLR results in a smoother and more stable descent in test loss, especially during the late training stages. This stable convergence leads to slightly better final performance compared to the traditional multi-step decay.
For the ViT-based Swin-Tiny model, we evaluate both CosineAnnealingLR and a Warmup + Cosine schedule. We find that adding a warm-up stage significantly enhances convergence speed and test loss. This is consistent with the behavior of transformer-based architectures, which are more sensitive to large initial learning rates. The Multi-step scheduler is not included in this setting, as it is generally less effective for transformers.
(a) ResNet-50: Multi-step vs. CosineAnnealingLR
(b) Swin-Tiny: CosineAnnealingLR vs. Warmup + Cosine
Figure 4: Log-scale test loss curves of KOALA++ on CIFAR-100 under different learning rate schedulers. Left: CosineAnnealingLR provides a smoother convergence than Multi-step on ResNet-50. Right: Warmup + Cosine improves stability and convergence for Swin-Tiny compared to CosineAnnealingLR alone.
Note: We observed that the Swin-Tiny model tends to overfit in the later stages when trained for 200 epochs. While we report the final Top-1 and Top-5 error at epoch 200 in Table 7 for consistency, the plotted curves in Figure 4 only show the first 100 epochs. This decision was made to focus on the informative convergence behavior and improve the visual clarity of the comparison.
C.3 Different Initializations of KOALA++
The initialization of the directional covariance product $v_{1}$ in KOALA++ depends directly on the initial covariance matrix $P_{0}$. To study the impact of $\sigma$ on the performance of the model, we conducted an ablation experiment in which we vary $\sigma\in\{0.05,0.1,0.2\}$ while keeping other hyperparameters fixed.
We perform this experiment on the CIFAR-100 dataset using the ResNet-50 architecture, training for 200 epochs. The learning rate is set to $2.0$, and the Kalman process noise parameter $q$ is fixed at $0.2$. The results are summarized in Table 8.
Table 8: Effect of varying $\sigma$ on KOALA++ performance. Dataset: CIFAR-100, Model: ResNet-50, Epochs: 200, $q=0.2$, lr$=2.0$
$\sigma$ Top-1 Error (%) Top-5 Error (%)
0.05 22.95 6.60
0.10 22.04 6.11
0.20 21.80 5.89
We observe that increasing $\sigma$ leads to improved performance, suggesting that initializing $v_{1}$ with stronger magnitude provides better early-stage directionality, which helps escape suboptimal regions in the optimization landscape. The result is shown in Figure 5
(a) Log-scale Test Loss
(b) Top-1 Validation Error (%)
Figure 5: Effect of varying $\sigma$ on KOALA++ performance over 200 epochs on CIFAR-100. Larger $\sigma$ values yield more stable convergence and lower final error.
Appendix D Experiments Details
D.1 Hardware
All experiments reported in this work were conducted on a server equipped with a single NVIDIA H100 GPU with 80 gigabytes of VRAM and 128 gigabytes of RAM. Unless otherwise stated, all training and evaluation tasks were executed using this configuration.
D.2 Image Classification
D.2.1 HP Settings for CIFAR10/100
We evaluate KOALA++ on both CIFAR-10 and CIFAR-100 using a variety of architectures, including CNN-based models (e.g., ResNet [8]) and vision transformers (e.g., Swin-Tiny [16]). For CIFAR-10, we initialize the process noise $\sigma$ and uncertainty level $q$ to 0.1. The initial learning rate is set to 1.0. In contrast, for CIFAR-100, given the larger number of classes and relatively richer sample distribution, we adopt a more aggressive hyperparameter configuration: both $\sigma$ and $q$ are set to 0.2, and the initial learning rate is increased to 2.0. For both datasets, we apply a weight decay of $5\times 10^{-4}$ to all ResNet and other convolutional models.
When applying KOALA++ to Swin-Tiny, we perform a coarse grid search over the weight decay parameter, evaluating the values $\{1\times 10^{-1},1\times 10^{-2},5\times 10^{-3},1\times 10^{-4},5\times 10^{-4}\}$. Interestingly, we find that KOALA++ tends to perform best with a smaller weight decay value ($1\times 10^{-4}$), whereas other optimizers usually favor a higher decay value (e.g., $1\times 10^{-2}$).
We hypothesize that this is due to KOALA++’s implicit regularization behavior stemming from its dynamic covariance updates, which already exert a stabilizing effect on parameter updates. Therefore, combining KOALA++ with a strong explicit regularizer such as large weight decay may lead to underfitting, particularly in transformer-based models where overparameterization is more common. Hence, a smaller weight decay appears more suitable when using KOALA++ in this context.
D.2.2 More Results for CIFAR10/100
In the main body of the paper, we only visualized 100-epoch results on CIFAR-100 using ResNet-50 to compare KOALA++ with several baselines. Here, we provide more comprehensive results by extending the training to 200 epochs and including CIFAR-10 as well.
CIFAR10
Figure 6(Left) presents the Top-1 validation error on CIFAR-10 with ResNet-50 across 200 training epochs. To highlight the behavior in the final stage, we also plot a zoomed-in view of the validation error during epochs 160–200 in Figure 6(Right).
(a) Validation Error (Epoch 1–200)
(b) Validation Error (Epoch 160–200)
Figure 6: Validation Top-1 Error (%) on CIFAR-10 using ResNet-50. Left: Full 200-epoch curve. Right: Zoomed view of epochs 160–200.
Note:This figure shows one random seed; KOALA++ outperforms SGD on average across three runs.
CIFAR100
We additionally report the 200-epoch results on CIFAR-100 using ResNet-50, comparing KOALA++ with several standard optimizers. The log-scale test loss and Top-1 validation error are shown in Figure 7.
(a) Log-scale Test Loss
(b) Top-1 Validation Error (%)
Figure 7: Results on CIFAR-100 with ResNet-50 over 200 epochs. Left: Log-scale test loss. Right: Top-1 validation error.
D.2.3 Experiments for ImageNet32
We train both ResNet18 and ResNet50 models on the ImageNet32 dataset. Following standard augmentation practices for low-resolution image classification, we apply RandomCrop(32, padding=4) and RandomHorizontalFlip() during training. The input images are subsequently converted to tensors and normalized using the standard ImageNet statistics. For validation, we directly normalize the images without applying any cropping or resizing. All models are trained for 100 epochs with a learning rate scheduler that decays the learning rate by a factor of 0.1 at epochs 30, 60, and 90, and a weight decay of 0.0001.
Note: This setup differs slightly from the data augmentation strategy used in KOALA’s ImageNet experiments [3], where validation images are center-cropped prior to evaluation.
Figure 8 shows the log-scale test loss and Top-1 validation error for ResNet18. We observe that SGD reaches a lower test error faster in early training but tends to overfit slightly in the later stages. In contrast, KOALA++ shows improved generalization in the long run, despite being less stable early on.
Table 9 summarizes the final Top-1 and Top-5 errors for both ResNet18 and ResNet50. The results show that KOALA++ achieves comparable or better performance to SGD without requiring any task-specific hyperparameter tuning.
Table 9: Final Top-1 and Top-5 validation error (%) on ImageNet32.
Model Optimizer Top-1 Error Top-5 Error
ResNet18 SGD 45.81 22.02
KOALA++ 45.71 21.94
ResNet50 SGD 40.17 17.57
KOALA++ 39.03 17.12
Figure 8: Left: Log-scale test loss; Right: Top-1 validation error for ResNet18 on ImageNet32.
D.3 Language Modeling
We adopt the experimental benchmark setup provided by AdaFisher [6], available at https://github.com/AtlasAnalyticsLab/AdaFisher. For a fair comparison, we use the same experimental pipeline and only test on the WikiText-2 dataset using the small GPT1 model.
To tune the hyperparameters of KOALA++, we performed a grid search over the following ranges:
  • Learning Rate (lr): {1e-1, 5e-2, 2e-2, 1e-2, 5e-3, 2e-3, 1e-3, 5e-4}
  • sigma and q (set equal): {0.01, 0.02, 0.05, 0.1}
  • Weight Decay: {1e-1, 5e-2, 2e-2, 1e-2, 5e-3, 2e-3, 1e-3, 5e-4, 2e-4, 1e-4}
The optimal configuration we found was: Learning Rate: 2e-3, sigma = q: 0.1, Weight Decay: 1e-4, which achieved the best perplexity among all tested settings. Figure 9 displays the validation loss and test perplexity curve of KOALA++ on the WikiText-2 dataset.
(a) Validation PPL vs Time (log)
(b) Validation Loss vs Time (log)
(c) Validation PPL vs Epoch (log)
Figure 9: Performance of KOALA++ on WikiText-2 with the small GPT1 model. Left: Log-scale validation perplexity over wall clock time. Right: Log-scale validation perplexity over training epochs. Following [6], AdamW was trained for 55 epochs, while KOALA++ were trained for 50 epochs.