### v2

# KOALA++: Efficient Kalman-Based Optimization with Gradient-Covariance Products

Abstract
We propose KOALA++, a scalable Kalman-based optimization algorithm that explicitly models structured gradient uncertainty in neural network training. Unlike second-order methods, which rely on expensive second order gradient calculation, our method directly estimates the parameter covariance matrix by recursively updating compact gradient covariance products. This design improves upon the original KOALA framework that assumed diagonal covariance by implicitly capturing richer uncertainty structure without storing the full covariance matrix and avoiding large matrix inversions. Across diverse tasks, including image classification and language modeling, KOALA++ achieves accuracy on par or better than state-of-the-art first- and second-order optimizers while maintaining the efficiency of first-order methods. The code is publicly available at https://github.com/Sumxiaa/KOALA_Plus_Plus.
1 Introduction
Optimization lies at the heart of deep learning. As neural networks continue to scale in depth, width, and data complexity, the efficiency, stability, and scalability of optimization algorithms have become increasingly important. First-order methods such as stochastic gradient descent (SGD) (robbins1951stochastic, ) and its adaptive variants like Adam (kingma2015adam, ) and RMSProp (tieleman2012lecture, ) offer practical efficiency and have driven much of the progress in deep learning. However, they often suffer from instability, slow convergence, and sensitivity to hyperparameter tuning – particularly in large-scale or noisy training regimes (keskar2017largebatch, ; zhang2019understanding, ).
To mitigate these issues, second-order optimizers such as K-FAC (martens2015optimizing, ) and Shampoo (gupta2018shampoo, ) introduce curvature-aware updates by approximating the Hessian or Fisher information matrix. These methods often improve convergence and generalization, but require costly matrix operations, limiting their scalability. Recently, AdaFisher (gomes2024adafisher, ) has emerged as a more scalable alternative, using block-diagonal Kronecker approximations of the Fisher information matrix to reduce overhead while preserving second-order benefits. Despite its effectiveness, AdaFisher fundamentally depends on the quality of the Fisher matrix estimation, which remains sensitive to noise and mini-batch variability.
An alternative and theoretically grounded direction comes from Kalman filtering. In Bayesian estimation theory, the posterior covariance matrix of an unbiased estimator defines the tightest achievable lower bound on parameter uncertainty, as formalized by the Cramér–Rao bound (rao1992information, ; cramer1999mathematical, ). Although the Fisher information matrix serves as an inverse bound proxy, it does not directly capture the true uncertainty in the parameters. Kalman filters, on the other hand, propagate the full covariance matrix of the estimate through time, providing a more faithful representation of uncertainty – particularly in noisy or nonstationary environments.
Motivated by this insight, KOALA (davtyan2022koala, ) proposed to frame stochastic optimization as a Kalman filtering problem, which models parameter updates as recursive state estimation, using a covariance matrix to adjust step sizes adaptively. However, KOALA’s original formulation relies on overly simplistic approximations, such as diagonal covariances and idealized gradient assumptions, which limit its practical effectiveness.
In this paper, we introduce KOALA++, a computationally efficient extension of the KOALA framework that explicitly maintains structured gradient covariance information. KOALA++ tracks the evolution of uncertainty directly through a recursive update at each training iteration $k$ of the gradient-covariance product $v_{k}=H_{k}P_{k-1}$, which we call surrogate, where $H_{k}$ is the gradient of the loss function and $P_{k-1}$ is the prior estimate of the parameter covariance. The key idea in our method is to approximate the recursive gradient-covariance update so that it never uses the covariance matrix $P_{k-1}$. By doing so, KOALA++ effectively avoids expensive matrix inversions and storing the entire covariance matrix $P_{k-1}$, while benefiting from the lifted constraints on its structure. Moreover, KOALA++ retains scalability akin to first-order methods, while achieving optimization dynamics and final performance competitive with state-of-the-art second-order methods.
Our main contributions are summarized as follows: We propose KOALA++, a Kalman-inspired optimizer that explicitly propagates structured gradient uncertainty via covariance modeling. We derive an efficient recursive update for the surrogate $v_{k}$ based on a minimum-norm solution to the covariance estimation problem. We implement KOALA++ with minimal overhead and validate its effectiveness across diverse architectures and tasks, including the ResNet family he2016deep , vision transformers liu2021swin , and language models. Extensive experiments show that KOALA++ consistently matches or exceeds the accuracy of first- and second-order baselines while offering improved stability and a better efficiency/performance trade-off.
2 Related Work and Motivation
2.1 Optimization Methods: A Brief Overview
First-order optimizers such as SGD robbins1951stochastic and Adam kingma2015adam are widely adopted for their scalability and simplicity, but often ignore the structure of gradient noise and require manually tuned learning rate schedules.
Second-order optimization methods, including K-FAC martens2015optimizing , Shampoo gupta2018shampoo , and more recently AdaFisher gomes2024adafisher , seek to address these shortcomings by explicitly incorporating curvature information. Nevertheless, practical implementations rely heavily on fixed approximations, such as block-diagonal or Kronecker-factored representations of the Hessian or Fisher matrices, to maintain computational feasibility. However, such approximations either still show limitations in terms of efficiency, when compared to first order methods, or adopt drastic simplifications that make the estimation process less stable.
2.2 Optimization via Kalman Filtering
The idea of using Kalman filtering for optimization has a long history. Early works include scalar Kalman updates for stochastic optimization ismail2018estimation , applications in reinforcement learning shashua2019trust , and connections between the Extended Kalman Filter (EKF) and natural gradient methods ollivier2019extended . Although these approaches demonstrate the versatility of Kalman-based updates, they are often task-specific or limited to small-scale models, and computing full covariance updates remains prohibitively expensive.
More recent research has revisited the Kalman filtering perspective in the context of large-scale neural optimization. Duran-Martin et al. duran2022efficient and Chang et al. chang2023low interpret learning as a Bayesian state estimation problem, where the network parameters are treated as latent states, and the likelihood of the data provides an explicit probabilistic observation model. These methods apply low-rank or subspace-structured versions of the Extended Kalman Filter to maintain tractable posterior updates. Their formulation operates in the full parameter or feature space (i.e., with matrix-valued state and observation models), which implies a computational cost of $\mathcal{O}(n^{2})$ to $\mathcal{O}(n^{3})$ per update, where $n$ is the number of model parameters, even under low-rank approximations. Consequently, they are mainly applicable to some relatively small scale settings.
In contrast, KOALA davtyan2022koala reinterprets gradient-based optimization as a scalar state estimation problem, where the loss function itself serves as an implicit observation. This design removes the need for an explicit probabilistic target and allows Kalman-style uncertainty propagation to be applied directly to general stochastic objectives. The formulation therefore scales linearly with the number of parameters, achieving $\mathcal{O}(n)$ complexity in practice, while remaining compatible with arbitrary differentiable loss functions. This implicit observation model makes KOALA more general and applicable to a wider range of optimization problems than Kalman filters formulated on explicit probabilistic observations. Nevertheless, the original KOALA formulation still involves certain simplifying assumptions that limit its expressiveness and stability in complex, anisotropic optimization landscapes.
For comparison, SLANG mishkin2018slang focuses purely on efficient covariance estimation through the diagonal plus low-rank structure without invoking any Kalman filtering or Bayesian state-update interpretation. Although SLANG achieves improved posterior approximations compared to diagonal methods, its updates are not dynamically informed by temporal prediction-correction as in Kalman filtering, and its computational cost remains at least $\mathcal{O}(n^{2})$.
In general, these methods differ in their dimensionality, observation modeling, and scalability. A gap remains between the expressiveness of structured covariance estimation and the efficiency of scalar Kalman filtering. This motivates the design of KOALA++, which preserves KOALA’s first-order scalability while introducing directionally aware covariance updates for large-scale optimization.
2.3 Motivation for KOALA++: Towards Structured Covariance Updates
The original KOALA formulation addresses efficiency by interpreting optimization as a scalar Kalman filtering problem, where the loss function acts as an implicit observation. However, its covariance update assumes $P_{k}\approx\sigma^{2}I$ and substitutes $H_{k}^{\top}H_{k}$ with gradient norms, which may reduce some information to a single scalar, limiting its ability to adapt to anisotropic curvature or correlated gradient directions, particularly in deep or noisy regimes.
On the other hand, as discussed in subsection 2.2, better approximations of the covariance matrix can provide richer uncertainty representations, but they are computationally prohibitive. This trade-off leaves a gap between expressiveness and scalability that current methods have yet to bridge.
Thus, KOALA++ is designed to overcome this limitation while retaining the first-order scalability of KOALA. Rather than maintaining or constraining a full covariance matrix, KOALA++ explicitly tracks a directional covariance projection
$v_{k}:=H_{k}P_{k-1},$
which quantifies how uncertainty evolves along current gradient directions. This projection is updated via an approximate recursive least-squares step that implicitly captures curvature anisotropy without requiring explicit matrix storage or inversion. As a result, KOALA++ provides a lightweight mechanism for propagating structured uncertainty, offering an effective balance between expressiveness and computational efficiency.
3 KOALA++
3.1 Problem Formulation and Kalman Filtering Notation
Neural network training is fundamentally an optimization problem, where, given a dataset ${\cal D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, the objective of the learning task is to adjust the model’s parameters $\theta\in\mathbb{R}^{n}$ so that the network’s predictions $f(x_{i};\theta)$ minimize the empirical risk
$\min_{\theta}{\cal L}(\theta),\quad\text{with }{\cal L}(\theta)=\frac{1}{N}\sum_{i=1}^{N}l(f(x_{i};\theta),y_{i}),$ (1)
where $l$ is a chosen loss function. However, in modern deep learning applications, $N$ is often large, which makes full-batch optimization computationally infeasible. As a result, practitioners rely on stochastic optimization methods that work on minibatches of data and minimize estimates of the loss
$L_{k}(\theta)=\frac{1}{m}\sum_{i\in B_{k}}l(f(x_{i};\theta),y_{i}),$ (2)
where $B_{k}$ is a set of indices in the $k$-th minibatch and $m=|B_{k}|$ is the cardinality of the minibatch. Although minibatch-based approaches such as SGD robbins1951stochastic and its adaptive variants are effective and scalable, they introduce noise due to the stochastic nature of loss and gradient estimates.
Table 1: EKF equations from davtyan2022koala .
$S_{k}=H_{k}(P_{k-1}+QI)H_{k}^{\top}+R$ innovation covariance, where $H_{k}=\nabla L_{k}^{\top}(\theta_{k-1})$
$K_{k}=(P_{k-1}+QI)H_{k}^{\top}S_{K}^{-1}$ Kalman gain
$P_{k}=(I-K_{k}H_{k})(P_{k-1}+QI)$ posterior of the state update
$\theta_{k}=\theta_{k-1}+K_{k}(L_{k}^{\text{target}}-L_{k}(\theta_{k-1}))$ state estimate update
To mitigate this issue, a recent approach, KOALA davtyan2022koala , casts neural network training as a recursive state estimation problem using Kalman filtering kalman1960new , a principled Bayesian framework for recursively estimating the hidden state of a system from noisy observations. At each iteration $k$, it combines prior estimates with new measurements to update the state $\theta_{k}$ and its associated uncertainty in the form of a covariance matrix $P_{k}$. Importantly, by keeping track of an estimate of the system’s state’s covariance, Kalman filtering is essentially implicitly using second order information about the state variables. The standard Extended Kalman Filter (EKF) generalizes this to non-linear systems using local linearization. In KOALA, model parameters $\theta_{k}$ are treated as latent states with stationary evolution under process noise, and the minibatch loss function $L_{k}$ serves as a noisy observation of a target loss value $L_{k}^{\text{target}}$. In these settings, the EKF equations can be summarized as in Table 1. To remain tractable in high-dimensional settings, KOALA simplifies the EKF update by:
Assuming an identity-scaled covariance matrix $P_{k}\approx\sigma_{k}^{2}I$; Approximating $H_{k}^{\top}H_{k}$ with $\|H_{k}\|^{2}I$; Reducing all matrix operations to scalar computations.
The resulting update closely resembles SGD, but with an adaptively scaled learning rate governed by the uncertainty estimate $P_{k}$.
While KOALA demonstrates promising stability and simplicity, its reliance on scalar approximations limits its capacity to capture directional or structured noise in high-dimensional loss landscapes. To address this, we propose KOALA++, which enhances the original framework with an efficient estimation of the directional state covariance while preserving the Kalman-based update algorithm.
3.2 Kalman Update with $v_{k}$-Based Reparameterization
To overcome the limitations of isotropic covariance assumptions in KOALA, KOALA++ introduces a more expressive formulation that captures directional uncertainty without incurring the full cost of matrix-valued updates. Instead of updating the full posterior covariance $P_{k}\in\mathbb{R}^{n\times n}$, we propagate a low-rank surrogate
$v_{k}:=H_{k}P_{k-1}\in\mathbb{R}^{1\times n},$ (3)
which represents the projection of the parameter covariance along the current measurement direction. This allows KOALA++ to maintain a richer understanding of uncertainty and curvature structure while keeping the computational complexity comparable to first-order methods.
Low-Rank Reparameterization of the Kalman Update:
By assuming a constant process noise $Q_{k}=Q$ for all $k$, the innovation covariance $S_{k}$ becomes
$S_{k}$ $=H_{k}(P_{k-1}+Q)H_{k}^{\top}+R=H_{k}v_{k}^{\top}+H_{k}QH_{k}^{\top}+R.$ (4)
Using this expression, the Kalman gain becomes
$K_{k}$ $=\frac{(P_{k-1}+Q)H_{k}^{\top}}{S_{k}}=\frac{v_{k}^{\top}+QH_{k}^{\top}}{S_{k}}$ (5)
and the corresponding state update is
$\theta_{k}\leftarrow\theta_{k}+\left(L_{k}^{\text{target}}-L_{k}(\theta_{k})\right)\frac{v_{k}^{\top}+QH_{k}^{\top}}{S_{k}}.$ (6)
Notice that, as suggested in davtyan2022koala , the learning rate of the update can be controlled by selecting a specific $L_{k}^{\text{target}}$. For example, $L_{k}^{\text{target}}=(1-\eta_{k})L_{k}(\theta_{k-1})$ ensures that the update is scaled by $\eta_{k}$.
Recursive Update of $v_{k}$:
To avoid storing $P_{k-1}$ explicitly, we recursively update $v_{k}$ by substituting the Kalman covariance update $P_{k-1}=(I-K_{k-1}H_{k-1})(P_{k-2}+Q)$
$v_{k}$ $=H_{k}P_{k-1}=H_{k}(I-K_{k-1}H_{k-1})(P_{k-2}+Q)$
$=\left(H_{k}-H_{k}\frac{v_{k-1}^{\top}+QH_{k-1}^{\top}}{S_{k-1}}H_{k-1}\right)(P_{k-2}+Q)$ (7)
$=H_{k}P_{k-2}+H_{k}Q-H_{k}\frac{v_{k-1}^{\top}+QH_{k-1}^{\top}}{S_{k-1}}v_{k-1}-H_{k}\frac{v_{k-1}^{\top}+QH_{k-1}^{\top}}{S_{k-1}}H_{k-1}Q.$
For simplicity, we define the recurring term
$\lambda_{k}$ $\doteq H_{k}\frac{v_{k-1}^{\top}+QH_{k-1}^{\top}}{S_{k-1}},$ (8)
which allows us to rewrite the update for $v_{k}$ in a more compact way:
$v_{k}=H_{k}P_{k-2}-\lambda_{k}v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q.$ (9)
In the last equation, all terms except the first can be efficiently calculated in iteration $k$. The first term cannot be easily evaluated due to the term $P_{k-2}$. The objective is to never store it explicitly to avoid the quadratic storage demand. As a first attempt, one could approximate $H_{k}$ with $H_{k-1}$ and replace the problematic term with $v_{k-1}$. However, we observed experimentally that this does not work well because of the high variability of the minibatch gradients.
As an alternative, we aim to retrospectively approximate the prior covariance matrix $P_{k-2}$ instead. We do this by taking advantage of the relation $v_{k-1}=H_{k-1}P_{k-2}$. However, since this equation is generally underdetermined in high-dimensional settings, we formulate a regularized inverse problem to obtain a unique and tractable solution. Specifically, we adopt the minimum Frobenius-norm formulation
$\min_{P_{k-2}}\|P_{k-2}\|_{F}^{2}\quad\text{subject to}\quad H_{k-1}P_{k-2}=v_{k-1},$ (10)
which yields a closed-form least-squares solution. This construction allows us to approximate $P_{k-2}$ efficiently without explicitly storing or inverting full covariance matrices.
For computational and memory efficiency, we use only one linear constraint, $v_{k-1}=H_{k-1}P_{k-2}$, although more could be used in principle. However, introducing more constraints, e.g., projections of other previous gradients $H_{k-2},H_{k-3}$, etc., would also require memorizing more gradient terms and using more calculations to update $v_{k}$. The use of the minimum Frobenius-norm criterion is to ensure a stable update for $v_{k}$ by limiting the norm of $P_{k-2}$.
In the next subsection, we discuss two variants of our algorithm, depending on the nature of the solution to the optimization problem (10).
3.3 Least-Squares Estimate of $P_{k-2}$
Variant I: Vanilla Least-Squares Estimation
In the first variant of our approach, we solve the least-squares problem without imposing any structural constraints on the covariance estimate. Given the relation $v_{k-1}=H_{k-1}P_{k-2}$, we compute $P_{k-2}\in\mathbb{R}^{n\times n}$ by introducing the Lagrange multipliers $\mu\in\mathbb{R}^{1\times n}$ and write the Lagrangian
$\mathcal{L}(P_{k-2},\lambda)=\|P_{k-2}\|_{F}^{2}+(H_{k-1}P_{k-2}-v_{k-1})\mu^{\top}.$ (11)
This formulation admits the closed-form solution [^1] See Appendix B.1 for a detailed derivation.
$P_{k-2}=\frac{H_{k-1}^{\top}v_{k-1}}{\|H_{k-1}\|^{2}}.$ (12)
By substituting this solution into $H_{k}P_{k-2}$ gives
$H_{k}P_{k-2}=\alpha_{k}v_{k-1},\quad\text{where}\quad\alpha_{k}:=\frac{H_{k}H_{k-1}^{\top}}{\|H_{k-1}\|^{2}}.$ (13)
Finally, we can obtain the following vanilla $v_{k}$ update
$v_{k}=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q.$ (14)
Variant II: Symmetric Covariance Estimation
To improve stability and theoretical consistency with classical Kalman filtering, we introduce a second variant that enforces symmetry in the approximation of $P_{k-2}$. In particular, the additional regularization term $P_{k-2}=P_{k-2}^{\top}$ is added to the objective. As in the previous derivation, we obtain a closed-form solution for $P_{k-2}$ [^2] See Appendix B.2 for a detailed derivation.
$P_{k-2}=\frac{H_{k-1}^{\top}\,v_{k-1}+v_{k-1}^{\top}\,H_{k-1}}{\|H_{k-1}\|^{2}}-H_{k-1}\,v_{k-1}^{\top}\,\frac{H_{k-1}^{\top}\,H_{k-1}}{\|H_{k-1}\|^{4}}.$ (15)
When substituting this into $H_{k}P_{k-2}$, the only difference between the two variants is the term
$r_{k}$ $=\frac{(H_{k}v_{k-1}^{\top})(H_{k-1}H_{k-1}^{\top})-(H_{k-1}v_{k-1}^{\top})(H_{k}H_{k-1}^{\top})}{\|H_{k-1}\|^{4}}.$ (16)
To simplify notation and implementation, we adopt a unified update rule for both variants. In the unconstrained case, we simply set $r_{k}=0$, effectively reducing the update to the asymmetric version. This allows the general form of the update to be written compactly as
$\boxed{v_{k}=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q+r_{k}H_{k-1}.}$ (17)
To maintain consistency with the original KOALA formulation and to simplify the calculation in the early stages of training, we initialize the directional covariance product $v_{1}=H_{1}P_{0}$ using a scaled identity assumption on the initial covariance matrix $P_{0}$. Specifically, we set:
$v_{1}:=\sigma_{0}^{2}H_{1},$ (18)
where $\sigma_{0}^{2}$ is a scalar hyperparameter representing the initial uncertainty. This corresponds to assuming $P_{0}=\sigma_{0}^{2}I$, which aligns with the isotropic prior used in KOALA and allows us to avoid explicit matrix computations at initialization. The general procedure of KOALA++ is summarized in Algorithm 1.
Algorithm 1 KOALA++
 Initialize $\theta_{0}$, $v_{1}$, $Q$, $R$, and fix the learning rate schedule $\eta_{k}$
for $k=2$ to $T$ do
  For simplicity, denote $H_{k}=\nabla L_{k}^{\top}(\theta_{k-1})$Calculate $\alpha_{k},\lambda_{k},r_{k}$ respectively from Equations (13), (8), and (16)
  Update:
$v_{k}$ $=(\alpha_{k}-\lambda_{k})v_{k-1}+(H_{k}-\lambda_{k}H_{k-1})Q+r_{k}H_{k-1}$ (19)
$\theta_{k}$ $=\theta_{k-1}-\frac{\eta_{k}L_{k}(\theta_{k-1})}{H_{k}v_{k}^{\top}+H_{k}QH_{k}^{\top}+R}\cdot(v_{k}^{\top}+QH_{k}^{\top})$ (20)
end for
return $\theta_{T}$
4 Experiments
We evaluate KOALA++ on a range of vision and language tasks to demonstrate its generality, stability, and convergence behavior compared to existing optimizers. Our experiments are organized into three parts: image classification, language modeling and ablation studies. Due to space constraints, we report a representative subset of results in the main text. Additional evaluations, including ablations, hyperparameter tuning, and extended visualizations, are provided in Appendix E.
4.1 Image Classification
CIFAR-10/100 Classification:
We follow the experimental setup of the original KOALA paper davtyan2022koala for CIFAR-10 and CIFAR-100 classification tasks, including data augmentation, model architectures, and optimization settings. We introduce two modifications for CIFAR-10: (i) training is repeated with three random seeds (42, 3407, and 2025) to ensure statistical robustness, and (ii) we adopt a two-stage learning rate decay strategy at the 100th and 150th epochs for 200-epoch training (inspired by successful multi-step decay strategies reported in prior works such as rice2020overfitting ), which improves generalization compared to the single-step decay used in the original paper. The results [^3] Some of the earlier reported results were obtained from existing public benchmarks, and therefore mean and standard deviation values are not available for all entries. To ensure reproducibility, we reran the experiments under identical settings; the full code and configurations are available at: https://github.com/Sumxiaa/KOALA_Plus_Plus. are shown in Table  2.
Table 2: Experimental results on CIFAR-10 and CIFAR-100 with different optimizers for 100 and 200 epochs. Best (blue) and second best (orange) performances are in boldface and with shading.
Dataset Architecture Method Error
100-epochs 200-epochs
Top-1 Err Top-5 Err Top-1 Err Top-5 Err
[…]
Adam 6.960.06 0.270.05 6.350.13 0.270.04
KOALA-M 6.290.04 0.370.02 6.060.14 0.280.06
[…]
KOALA-M 8.090.16 0.290.04 7.780.39 0.320.06
[…]
KOALA-M 7.620.13 0.300.04 7.670.43 0.320.06
[…]
Adam 24.250.34 7.060.15 24.140.12 7.470.34
[…]
W-ResNet-50-2 SGD 23.030.29 5.970.11 23.500.15 6.470.16
Adam 23.990.17 6.690.07 23.740.19 7.100.17
[…]
Comparison with Recent Advances
To comprehensively evaluate the effectiveness and generalization capability of KOALA++, we design two experimental settings that progressively incorporate more recent optimizers, modern network architectures, stronger learning rate schedulers, and richer data augmentations. This design mirrors recent benchmarking practices in the literature, particularly those from KOALA davtyan2022koala and AdaFisher gomes2024adafisher .
(1) CIFAR-100 + ResNet-50 + Advanced Optimizers. To evaluate KOALA++ in a standardized yet competitive setting, we follow the benchmarking protocol introduced in KOALA davtyan2022koala , and compare our method with a wide range of recent optimizers on the CIFAR-100 dataset using ResNet-50, trained for 100 epochs. All optimizers are tested under the same learning rate schedule (StepLR) and training configuration. For each optimizer, we adopt hyperparameters as suggested in the original papers or official implementations. The settings and results are shown in Table 3.
Table 3: CIFAR-100 (ResNet-50, 100 epochs, StepLR) results and optimizer hyperparameters.
Optimizer Top-1 Err. Top-5 Err. Hyperparameters / Source
[…]
Adamax kingma2015adam 32.42 10.74 Same as Adam
AdamW loshchilov2017decoupled 27.23 7.98 Same as Adam
AdamP heo2020adamp 26.62 7.61 Same as Adam
Amsgrad reddi2019convergence 25.27 6.78 Same as Adam
Adan xie2024adan 24.92 6.86 Paper and repo defaults [^4] https://github.com/lucidrains/Adan-pytorch
[…]
AdaFisher gomes2024adafisher 23.38 6.05 Original AdaFisher hyperparams
Adabelief zhuang2020adabelief 23.07 6.05 Official config [^6] https://github.com/juntang-zhuang/Adabelief-Optimizer#
KOALA-M 23.71 6.58 Redid the experiments with our implementation settings
[…]
(a) Test Loss (log)
(b) Test Top-1 Error
Figure 1: Comparison of training loss, validation loss, and validation error for different optimizers on CIFAR-100 using ResNet-50. KOALA++ demonstrates the strongest performance drop at scheduled learning rate decays (epochs 30, 60, and 90), highlighting its superior scheduler responsiveness.
Figure 1 presents the training and validation dynamics on CIFAR-100 with ResNet-50 under a multi-step learning rate schedule at epochs 30, 60, and 90. Compared to SGD, Adam, AdaFisher, and KOALA-M, our proposed KOALA++ demonstrates consistently lower validation loss on the logarithmic scale, along with improved generalization as shown by the lower Top-1 error.
In particular, the benefits of KOALA++ are most evident as the scheduled learning rate drops: the optimizer exhibits sharp improvements in both loss and error immediately following the milestones, indicating more effective adaptation to the changing optimization dynamics. This suggests that KOALA++ better integrates structural information from past gradients and uncertainty estimates, allowing it to respond more effectively to scheduler-induced shifts.
Moreover, KOALA++ maintains competitive or superior stability throughout all training phases, without exhibiting the overfitting spikes or stagnation observed in other optimizers. These observations collectively validate KOALA++ as a robust and efficient optimizer for large-scale vision tasks.
(2) CIFAR-10 + Modern Models + Cosine Annealing + Strong Augmentation To align with the stronger benchmark protocol proposed by AdaFisher gomes2024adafisher , we conduct a second set of experiments that adopt more modern practices in deep learning optimization:
  • Advanced models: ResNet-101, DenseNet-121 huang2017densely , MobileNetV3 howard2019searching , and the Swin Transformer (Tiny) liu2021swin , reflecting a diversity of architectures including convolutional and attention-based designs.
  • Modern learning rate scheduler: We use CosineAnnealingLR for all optimizers, which has been shown to improve convergence stability and final accuracy in modern setups.
  • Stronger data augmentation: In addition to standard cropping and flipping, we apply Cutout devries2017improved during training to improve generalization.
  • Larger batch size: Following current best practices, we increase the batch size to 256, which can benefit optimizers that are sensitive to noisy gradients.
  • More Random Seeds: Following the AdaFisher benchmark protocol gomes2024adafisher , we report top-1 accuracy averaged over five random seeds: 42, 3407, 9331, 7, and 2025. The standard deviation across these runs is shown in the bottom-right corner of each table entry.
Table 4: Top-1 accuracy (%) on CIFAR-10 across various architectures. We report the average over 5 seeds. The large number is the mean and the subscript is the standard deviation.
Optimizer ResNet50 ResNet101 DenseNet121 MobileNetV3 Tiny Swin
SGD 95.710.1 95.980.2 96.090.1 94.430.2 82.340.2
Adam 94.450.2 94.570.1 94.860.1 93.320.1 87.370.6
AdaHessian 95.540.1 95.290.6 96.110.1 92.860.1 84.150.2
K-FAC 95.660.1 96.010.1 96.120.1 94.340.1 64.790.5
Shampoo 94.590.1 94.630.1 95.660.1 93.810.2 63.910.4
[…]
As shown in Table 4, KOALA++ performs on par with AdaFisher across a range of architectures, while exhibiting consistently lower standard deviations, indicating greater stability among random seeds. All KOALA++ results were obtained without extensive hyperparameter adjustment, highlighting the robustness of the method in the default settings. It is also worth noting that on Swin-Tiny, KOALA++ reaches strong performance with significantly smaller variance, demonstrating better adaptability to Transformer-based architectures.
Efficiency profiling.
Beyond accuracy, we also compare the computational efficiency of KOALA++ with strong baselines. KOALA++ achieves runtime and memory comparable to Adam(W), while being substantially more efficient than Shampoo and AdaHessian. A detailed comparison, including per-epoch time, FLOPs, and peak memory usage on ResNet-50 and Swin-Tiny, is provided in the Appendix E.3.
4.2 Language Model
We evaluated KOALA++ on the Wikitext-2 dataset, following the same language modeling benchmark setting introduced in the AdaFisher paper gomes2024adafisher . The model is a scaled-down version of GPT-1 with four masked self-attention layers and approximately 28 million parameters. We adopt the same architecture, tokenizer, and 50-epoch training schedule as used in the original benchmark.
In line with the original AdaFisher benchmark (see Github [^7] https://github.com/AtlasAnalyticsLab/AdaFisher/tree/main), we report the test perplexity (PPL) corresponding to the best-performing model on the validation set, along with the total training time, to provide a holistic view of performance and efficiency. The results are summarized in Table 5. In addition to reporting test perplexity and total training time, we further provide visualizations to better illustrate the training dynamics (see Appendix E).
4.3 Ablation Study
Due to space constraints, we report only the core ablation comparing KOALA-M and two variants of KOALA++ on CIFAR-100 with ResNet-50. The default KOALA++ uses a symmetric gain approximation, while KOALA++ (NS) adopts an asymmetric covariance. As shown in Figure 2, KOALA++ achieves faster convergence and better final performance. KOALA++ (NS) is competitive but less stable in later stages, underscoring the benefit of structural symmetry. Full ablations are provided in the Appendix D.
(a) Test Loss (log)
(b) Test Top-1 Error
Figure 2: Ablation study comparing KOALA++, its asymmetric variant KOALA++ (NS), and KOALA-M on CIFAR-100 with ResNet-50. Left: Log-scale test loss curves. Right: Validation Top-1 error rates.
Table 5: Language modeling performance on Wikitext-2. We report the test perplexity (lower is better) corresponding to the best validation model, as well as the total training time over 50 epochs.
Optimizer Test PPL ↓ Average Training Time per Epoch (s)
AdamW 173.98 23.11
AdaHessian 407.69 62.91
[…]
5 Conclusions, Limitations and Future Work
We proposed KOALA++, an efficient optimization method that extends the Kalman-based KOALA algorithm by incorporating a lightweight directional covariance approximation. KOALA++ retains the stability and simplicity of its predecessor, while substantially improving its capacity to capture general structured uncertainty in high-dimensional loss landscapes. Through extensive experiments across vision and language tasks, we demonstrate that KOALA++ achieves competitive or superior performance compared to both first- and second-order optimizers, while maintaining training efficiency on par with standard first-order methods.
Limitations. A notable limitation of our method is that the approximated covariance matrix $P_{k}$ is not explicitly constrained to be positive semi-definite (PSD). Although we empirically observe that the eigenvalues of $P_{k}$ remain stable and bounded throughout the training (see the Appendix B.3), the lack of a formal PSD guarantee may pose risks in certain settings or applications involving sensitive numerical stability. Addressing this limitation efficiently is left to future work.
Future Work. To address this issue, future work may explore enforcing PSD constraints via projection, spectral regularization, or reparameterization techniques. Additionally, we plan to investigate the integration of KOALA++ with adaptive gradient clipping and apply it to large-scale pretraining scenarios such as vision-language models, diffusion models, or instruction-tuned LLMs.
Acknowledgements
Calculations were performed on UBELIX (https://www.id.unibe.ch/hpc), the HPC cluster at the University of Bern. Aram Davtyan has been supported by SNSF Grant 10001278.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state our main contributions and support them with theoretical and experimental results (see Sections 1 , 3, and 4).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations in Section  5, including assumptions on covariance symmetry and potential instability.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The assumptions are made in Section  3 and the derivations are provided in the Appendix B.1 B.2
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All experimental settings are described in Section 4 and Appendix E, including models, datasets, and hyperparameters.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We plan to release the code upon publication, and all datasets used (CIFAR-10/100, WikiText-2) are publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Details of model architectures, optimizer configurations, learning rates, and schedules are provided in Section  4 and Appendix E.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report averages over 3–5 seeds and include standard deviations in Table 4 and Appendix D.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe compute resources in Appendix E, including training time on an H100 GPU.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We adhere to the NeurIPS Code of Ethics. No private or sensitive data is used.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: The paper does not discuss societal impacts, as it focuses solely on the design and evaluation of optimization algorithms, which are technical contributions without direct societal applications.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not describe such safeguards, as it does not involve the release of high-risk models or datasets such as pretrained language models, image generators, or scraped data.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All external assets used in the paper (e.g., the AdaFisher benchmark) are properly credited, and their licenses and terms of use (e.g., MIT License) have been respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The reimplementation of the KOALA++ optimizer is a new asset introduced in this paper. It is well documented, and usage instructions will be provided alongside the code repository.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The study does not involve human participants, so IRB approval and risk disclosures are not applicable.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We used LLMs for proofreading and structuring (e.g., ChatGPT), but they were not used as part of the core methodology in this research.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix Contents
Appendix A KOALA
In this section, we are going to revisit the original KOALA framework, which was only briefly introduced in the main text. KOALA [4] is a Kalman-inspired optimization algorithm that interprets gradient-based updates as a recursive state estimation process. To ensure scalability in high-dimensional deep learning settings, the original work proposed two tractable variants: KOALA-V and KOALA-M. These variants differ in how they approximate the parameter covariance and in whether they incorporate momentum dynamics. In what follows, we provide a detailed and self-contained exposition of both KOALA-V and KOALA-M, including their derivations, update rules, and practical considerations. This not only offers a clearer understanding of KOALA as a foundational method, but also sets the stage for the improvements introduced in KOALA++.
A.1 Risk Minimization as Loss Adaptivity
[…]
${L}_{k}(\theta)\simeq{\cal L}(\theta)-v_{k},\quad\text{where }v_{k}\sim\mathcal{N}(0,R_{k}),$ (21)
$\theta$ denotes the model parameters and $R_{k}$ is the variance of the observation noise.
Rather than optimizing $\min_{\theta}{\cal L}(\theta)$ directly, KOALA reformulates the problem by adapting the model parameters so that the stochastic risk ${L}_{k}(\theta_{k})$ matches a desired target risk ${L}^{\text{target}}_{k}$, i.e.,
${L}_{k}(\theta_{k})={L}^{\text{target}}_{k}-v_{k}.$ (22)
This formulation enables the use of Kalman filtering, which provides a principled way to recursively update the estimate of $\theta_{k}$ by combining prior knowledge with new and noisy observations. The underlying system can be described as a standard dynamical model:
$\theta_{k}$ $=\theta_{k-1}+w_{k-1},\quad w_{k}\sim\mathcal{N}(0,Q_{k})$ (23)
${L}^{\text{target}}_{k}$ $={L}_{k}(\theta_{k})+v_{k},\quad v_{k}\sim\mathcal{N}(0,R_{k}).$ (24)
A.2 KOALA-V (Vanilla)
The idea behind KOALA-V is remarkably natural, which is to directly apply the standard Kalman filtering update equations to the system in Equations (23) and (24). However, directly computing the posterior covariance $P_{k}\in\mathbb{R}^{n\times n}$ is infeasible for high-dimensional models. To address this, KOALA-V introduces the following approximations:
  • The posterior covariance $P_{k}$ is approximated by a scaled identity matrix:
    $P_{k}\approx\sigma_{k}^{2}I_{n}.$ (25)
  • $H_{k}^{\top}H_{k}$ is approximated by the squared norm of the gradient:
    $H_{k}^{\top}H_{k}\approx\left\|\nabla_{\theta}{L}_{k}(\theta_{k})\right\|^{2}I_{n}.$ (26)
The resulting parameter update rule is given in Algorithm 2.
This update closely resembles the SGD update:
$\theta_{k}=\theta_{k-1}-\eta\nabla\hat{L}_{k}(\theta_{k}),$ (27)
but with a key difference: in KOALA-V, the scaling of the gradient is automatically adapted based on the current loss, gradient norm, and covariance estimate. This adaptivity makes KOALA-V more robust to noisy training dynamics than static first-order methods like SGD or Adam.
Algorithm 2 KOALA-V (Vanilla)
 Initialize $\theta_{0}$, $P_{0}$, $Q$ and $R$
for $k$ in range(1, $T$) do
  Predict:
$\hat{\theta}_{k}=\theta_{k-1};\quad\hat{P}_{k}=P_{k-1}+Q$
  Update:
$\theta_{k}$ $=\hat{\theta}_{k}-\frac{\hat{P}_{k}(\hat{L}_{k}(\hat{\theta}_{k})-\hat{L}_{k}^{\text{target}})}{\hat{P}_{k}\|\nabla\hat{L}_{k}(\hat{\theta}_{k})\|^{2}+R}\cdot\nabla\hat{L}_{k}(\hat{\theta}_{k})$ (28)
$P_{k}$ $=\frac{R}{\hat{P}_{k}\|\nabla\hat{L}_{k}(\hat{\theta}_{k})\|^{2}+R}\cdot\hat{P}_{k}$ (29)
end for
return $\theta_{T}$
A.3 KOALA-M (Incorporating Momentum Dynamics)
KOALA-M extends the original KOALA framework by explicitly integrating momentum-based dynamics into the Kalman filtering formulation. Instead of modeling the parameter update as a simple random walk, KOALA-M introduces an auxiliary momentum state $p_{k}$, resulting in the augmented state dynamics:
$\theta_{k}$ $=\theta_{k-1}+p_{k-1}+w_{k-1},\quad w_{k-1}\sim\mathcal{N}(0,Q),$ (30)
$p_{k}$ $=\kappa\,p_{k-1}+u_{k-1},\quad\quad\quad\quad u_{k-1}\sim\mathcal{N}(0,U),$ (31)
where $\kappa$ controls the momentum persistence. This momentum augmentation enables KOALA-M to better capture and leverage correlations across sequential updates, effectively enhancing stability.
However, KOALA-M inherits computational complexity due to its augmented state. To mitigate this, KOALA-M approximates the posterior covariance with a structured form, typically assuming it to be a Kronecker product of a $2\times 2$ covariance matrix with an identity matrix $I_{n}$. Similar to KOALA-V, KOALA-M employs layer-wise covariance updates to obtain a better approximation of parameter uncertainty, partially alleviating the limitation of overly simplistic covariance structures.
In contrast, our proposed KOALA++ algorithm does not require explicit momentum augmentation or restrictive covariance approximations due to its inherently richer and adaptive covariance representation. KOALA++ directly maintains structured gradient-covariance products without imposing additional structural assumptions on the covariance matrix $P$, thus naturally capturing both directional information and momentum-like dynamics implicitly.
Appendix B Mathematical Foundations of KOALA++
In the main paper, we provided both the non-symmetric and symmetric solutions for $P_{k-2}$. In this section, we provide the detailed derivations of those. For notational simplicity, where possible, we drop the subscripts in $k$, and we can restate the optimization problem as follows:
$\min_{P}\|P\|_{F}^{2}\quad\text{subject to}\quad HP=v$ (32)
B.1 Vanilla Optimal Solution of Covariance Matrix P
To solve this constrained optimization problem, we introduce a Lagrange multiplier $\mu\in\mathbb{R}^{1\times n}$ and construct the corresponding Lagrangian function:
$\mathcal{L}(P,\mu)=\|P\|_{F}^{2}+\mu(HP-v)^{\top}.$ (33)
We then find the optimal solution by differentiating the Lagrangian $\mathcal{L}$ with respect to $P$, and set this derivative to zero to find critical points:
$\frac{\partial\mathcal{L}(P,\mu)}{\partial P}=2P+H^{\top}\mu=0\quad\Rightarrow\quad P=-\frac{1}{2}H^{\top}\mu.$ (34)
Now, substituting this optimality condition into the constraint equation $HP=v$, we have:
$H\left(-\frac{1}{2}H^{\top}\mu\right)$ $=v\Rightarrow\mu=-2\frac{v}{HH^{\top}}.$ (35)
Thus, the optimal solution for $P$ is obtained by substituting back the expression for $\mu$:
$\boxed{P=\frac{H^{\top}v}{HH^{\top}}}.$ (36)
This solution is referred to as the vanilla optimal solution (asymmetric) because it does not explicitly impose symmetry constraints on $P$. In the next subsection, we derive a symmetric solution variant that respects the symmetry requirement typically imposed on covariance matrices.
B.2 Symmetric Optimal Solution of Covariance Matrix P
For derivation of the symmetric solution, we use the same Lagrangian framework, but now explicitly incorporate the symmetry constraint into the formulation. The Lagrangian function, considering the symmetry constraint, remains the same as before (see Equation (33)).
We use the symmetric matrix derivative form from the Matrix Cookbook [23].Taking the derivative with respect to $P$ and setting it to zero yields
$\frac{\partial\mathcal{L}}{\partial P}(\text{Symmetric})=\left[\frac{\partial\mathcal{L}}{\partial P}\right]+\left[\frac{\partial\mathcal{L}}{\partial P}\right]^{\top}-\text{diag}\left[\frac{\partial\mathcal{L}}{\partial P}\right]=0$ (37)
By plugging in Equation (33), we have
$2P+\mu^{\top}H+2P^{\top}+H^{\top}\mu-\text{diag}(2P+\mu^{\top}H)=0$ (38)
Since $P$ is symmetric, we can rewrite Equation (38) as:
$4P+\mu^{\top}H+H^{\top}\mu-\text{diag}(2P+\mu^{\top}H)=0,$ (39)
which gives the following equations
$\begin{cases}4P+\mu^{\top}H+H^{\top}\mu=0&\text{(off diagonal)}\\ 2P+H^{\top}\mu=0&\text{(on diagonal).}\end{cases}$ (40)
We observe that the off-diagonal solution satisfies the on-diagonal equation on the diagonal. Therefore, we only need to solve the off-diagonal case. From the off-diagonal constraint $4P+\mu^{\top}H+H^{\top}\mu=0$, we have $P=-\frac{1}{4}\left(\mu^{\top}H+H^{\top}\mu\right)$. By substituting in $HP=v$, we obtain
$HP=-\frac{1}{4}(H\mu^{\top}H+HH^{\top}\mu)=v.$ (41)
We can solve $\mu$ as
$\mu=-4v(H^{\top}H+HH^{\top}I)^{-1},$ (42)
where $I$ denotes the identity matrix. To compute the inverse, we recall the Woodbury matrix identity [23]:
$(A+UV)^{-1}=A^{-1}-A^{-1}U(I+VA^{-1}U)^{-1}VA^{-1}$ (43)
By letting $A=HH^{\top}I$, $U=H^{\top}$, and $V=H$, we find the following expression
$(HH^{\top}I+H^{\top}H)^{-1}=\frac{1}{HH^{\top}}\left(I-H^{\top}\cdot\frac{1}{2}H\cdot\frac{1}{HH^{\top}}\right).$ (44)
This provides an expression for $\mu$ without matrix inverses, i.e.,
$\mu=-4v\frac{1}{HH^{\top}}\left(I-\frac{1}{2}\frac{H^{\top}H}{HH^{\top}}\right).$ (45)
Finally, we can plug this expression to have a closed-form optimal solution $P^{*}$, that is,
$P^{*}$ $=\frac{1}{HH^{\top}}\left(I-\frac{1}{2}\frac{H^{\top}H}{HH^{\top}}\right)v^{\top}H+H^{\top}v\frac{1}{HH^{\top}}\left(I-\frac{1}{2}\frac{H^{\top}H}{HH^{\top}}\right)$
$=\frac{v^{\top}H+H^{\top}v}{HH^{\top}}-\frac{Hv^{\top}}{(HH^{\top})^{2}}H^{\top}H.$ (46)
B.3 Eigenvalues Analysis of the Optimal Matrix $P^{*}$
Recall the closed-form expression for the optimal matrix
$P^{*}=\frac{1}{HH^{\top}}(H^{\top}v+v^{\top}H)-\frac{Hv^{\top}}{(HH^{\top})^{2}}H^{\top}H.$ (47)
To analyze the definiteness of $P^{*}$, we consider its eigenvalues. Let $w\in\mathbb{R}^{n\times 1}$ be an eigenvector of $P^{*}$ and $\lambda$ its corresponding eigenvector, then,
$\lambda w=P^{*}w.$ (48)
We can immediately see that $w$ lies in the span of $H^{\top}$ and $v^{\top}$, because the range of $P$ lies in their span. Then, we can write
$w=\alpha H^{\top}+\beta v^{\top}.$ (49)
By substituting it into the eigenvalue equation we obtain
$\lambda(\alpha H^{\top}+\beta v^{\top})=\frac{1}{HH^{\top}}(H^{\top}v+v^{\top}H)(\alpha H^{\top}+\beta v^{\top})-\frac{Hv^{\top}}{(HH^{\top})^{2}}H^{\top}H(\alpha H^{\top}+\beta v^{\top}).$ (50)
We now compute both terms. For simplicity, let us name the following three scalar values as
$x=HH^{\top},\quad y=Hv^{\top},\quad z=vv^{\top}.$ (51)
Then, we have
$P^{*}w$ $=\frac{1}{x}(H^{\top}v\alpha H^{\top}+H^{\top}v\beta v^{\top}+v^{\top}H\alpha H^{\top}+v^{\top}H\beta v^{\top})$ (52)
$-\frac{1}{x^{2}}(Hv^{\top}H^{\top}H\alpha H^{\top}+Hv^{\top}H^{\top}H\beta v^{\top})$ (53)
$=\frac{1}{x}(\alpha yH^{\top}+\beta zH^{\top}+\alpha xv^{\top}+\beta yv^{\top})-\frac{1}{x^{2}}\alpha yxH^{\top}-\frac{1}{x^{2}}\beta y^{2}H^{\top}$ (54)
$=\frac{\beta z}{x}H^{\top}+\alpha v^{\top}+\frac{\beta y}{x}v^{\top}-\frac{\beta y^{2}}{x^{2}}H^{\top}$ (55)
$=\frac{\beta(xz-y^{2})}{x^{2}}H^{\top}+\frac{\alpha x+\beta y}{x}v^{\top}.$ (56)
By matching both sides coefficient-wise, we derive a 2D eigenvalue system as follows
$\alpha\lambda$ $=\dfrac{\beta(xz-y^{2})}{x^{2}}$ (57)
$\beta\lambda$ $=\dfrac{\alpha x+\beta y}{x}.$ (58)
If $\alpha=0$, then from Equation (57) we know that $\beta=0$ or $xz-y^{2}=0$. However, $\beta=0$ implies that $w=\alpha H^{\top}+\beta v^{\top}=0$, which contradicts the assumption that $w$ is a non-zero eigenvector. On the other hand, by the Cauchy-Schwarz inequality, $xz-y^{2}=0$ implies that the two directions $H^{\top}$ and $v^{\top}$ become linearly dependent. In this case, the eigenvalues of $P^{*}$ are 0 and $a$, where $H^{\top}=av^{\top}$, due to the linear dependence. The linear dependence is, however, very unlikely in practice, as, in general $H$ and $v$ span a two-dimensional subspace.
Therefore, we very likely have $\alpha\neq 0$. In this case, the eigenvalue $\lambda$ must satisfy
$\lambda=\frac{\beta}{\alpha}\cdot\frac{xz-y^{2}}{x^{2}}.$ (59)
Now, we need to solve the following equation to examine the sign of $\frac{\beta}{\alpha}$
$(xz-y^{2})\left(\frac{\beta}{\alpha}\right)^{2}-xy\frac{\beta}{\alpha}-x^{2}=0.$ (60)
We get
$\frac{\beta}{\alpha}=\frac{xy\pm\sqrt{x^{2}y^{2}+4(xz-y^{2})x^{2}}}{2(xz-y^{2})}.$ (61)
Finally, the eigenvalues of the matrix $P$ are
$\lambda_{1,2}=\frac{xy\pm\sqrt{x^{2}y^{2}+4(xz-y^{2})x^{2}}}{2x^{2}}.$ (62)
Although the matrix $P$ obtained through least-squares minimization does not guarantee positive definiteness, we conducted an empirical analysis by visualizing the evolution of its eigenvalues at different training stages.
Specifically, we performed eigenvalue tracking on the ResNet-50 model trained on CIFAR-100 at 100 epochs. The learning rate was decayed at epochs 30, 60, and 90, which correspond to the early, middle, and late stages of training, respectively.
To better understand the behavior of the matrix $P$, we analyzed the distribution of its eigenvalues across all mini-batches at four critical epochs: 0, 30, 60, and 90. The results are shown in Figure 3. We found the cases for 60 and 90 epochs to give identical trends as for 30 epochs case. Thus, to avoid unnecessary repetition, we are not showing the plots for these cases.
At epoch 0 (Figure 3, left), the eigenvalues exhibit a wide dynamic range with both large and small eigenvalues increasing significantly over time. The large eigenvalue (orange) and small eigenvalue (blue) rise rapidly and display sharp oscillations, indicating an evolving optimization landscape in the early stage of training.
However, by epoch 30 (Figure 3, right), the eigenvalue trajectories stabilize, and the magnitude of fluctuations reduces dramatically. From epoch 30 onwards, the eigenvalue distributions become increasingly similar. In these later epochs, both the large and small eigenvalues remain relatively flat and nearly constant throughout training, reflecting a stabilized matrix structure and learning dynamics.
(a) Epoch 0
(b) Epoch 30
Figure 3: Evolution of the large and small eigenvalues of the matrix $P$ at epoch 0 and epoch 30 across all mini-batches.
Appendix C Convergence Behavior of KOALA++
C.1 On the Stability and Implicit Positive Semi-Definiteness of KOALA++
While the current formulation of KOALA++ does not explicitly enforce the positive semi-definite (PSD) constraint on the covariance-like matrix, we empirically observe that the optimizer remains numerically stable and convergent across diverse architectures and datasets. This suggests that, in practice, the lack of an explicit PSD guarantee does not adversely affect convergence behavior.
This is probably because, when we are calculating the eigenvalues in Equation (62), the quantities $\lambda_{1,2}$ are obtained from the least-squares solution of the auxiliary matrix $P_{k-2}$, which represents the optimal estimate under the minimal-norm fitting objective. In other words, the eigenvalues in Equation (62) correspond to an analytically fitted surrogate of the covariance update, rather than the actual covariance-like matrix used in the algorithmic recursion. In practice, the true update matrix—hereafter referred to as the ground truth matrix—is defined by integrating the solution from Equation (12) into (9), then we have:
$v_{k}$ $=H_{k}\left(\frac{H^{\top}v+v^{\top}H}{\|H\|^{2}}-\frac{(Hv^{\top})(H^{\top}H)}{\|H\|^{4}}\right.\left.+QI-\frac{(v^{\top}+QH^{\top})(v+QH)}{H(v^{\top}+QH^{\top})+R}\right).$ (63)
We call the ground-truth matrix $M_{k}$ and express it as follows
$M_{k}$ $=\frac{H^{\top}v+v^{\top}H}{\|H\|^{2}}-\frac{(Hv^{\top})(H^{\top}H)}{\|H\|^{4}}+QI-\frac{(v^{\top}+QH^{\top})(v+QH)}{H(v^{\top}+QH^{\top})+R}.$ (64)
Based on this formulation, we can further examine the directional definiteness of the update. Specifically, by evaluating the product $H_{k}v_{k}^{\top}$, we can infer whether the update direction induced by $M_{k}$ corresponds to a positive-definite step along the current gradient direction. In particular, when
$H_{k}v_{k}^{\top}=H_{k}M_{k}H_{k}^{\top}>0,$ (65)
the update matrix $M_{k}$ can be considered positive definite with respect to the local curvature defined by $H_{k}$. This provides a practical criterion for assessing the stability and consistency of the covariance update in KOALA++.
Empirically, we verify this property on three benchmark datasets: CIFAR-10, CIFAR-100, and ImageNet32. For each case, we monitor the angle between the curvature vector $H_{k}$ and the update direction $v_{k}$ throughout the training. As shown in Figure 4, the angle remains consistently acute ($90^{\circ}$) in all training stages, which implies that $H_{k}v_{k}^{\top}>0$, thus confirming the numerical positive semi-definiteness of $M_{k}$ and the empirical stability of the update rule.
(a) CIFAR-10
(b) CIFAR-100
(c) ImageNet32
Figure 4: Empirical verification of the positive semi-definiteness of $M_{k}$ in KOALA++. The plots show the evolution of the angle between the curvature vector $H_{k}$ and the update direction $v_{k}$ across training for three datasets. The consistently acute angles ($90^{\circ}$) indicate that $H_{k}$ and $v_{k}$ remain directionally aligned, confirming the effective PSD behavior of $M_{k}$ in practice.
However, from a theoretical standpoint, ensuring positive-semidefiniteness could further strengthen the stability guarantees of the method. In particular, one promising direction is to modify the optimization objective so that the covariance estimate minimizes its distance—for example, in the Frobenius norm—to a structured PSD matrix such as a scaled identity. Such a formulation naturally leads to a convex optimization problem with a unique and bounded solution, thereby eliminating ambiguity and improving theoretical soundness.
Although this PSD-enforced variant is beyond the current scope of our study, we view it as an important extension to future work aimed at establishing stronger convergence guarantees for KOALA++.
C.2 Convergence Characteristics of KOALA++ vs. Quasi-Second-Order Optimizers
To better understand how KOALA++ differs from quasi-second-order optimizers in convergence behavior, we conduct a controlled comparison on CIFAR-10 using a ResNet-50 backbone trained for 100 epochs under a cosine learning rate scheduler.
Although the Shampoo optimizer [8] is a well-established quasi-second-order method, its computational overhead makes direct comparison impractical under identical epoch budgets. To ensure a fair and efficient evaluation, we instead adopt Muon [14], a more recent and scalable curvature-aware optimizer that follows a similar quasi-second-order philosophy but operates with significantly lower wall-clock cost. This substitution allows us to better examine the convergence profile of KOALA++ against comparable quasi-second-order approaches.
(a) Log-scale Test Loss vs Epochs
(b) Top-1 Validation Error vs Epochs
Figure 5: CIFAR-10 with ResNet18 (Left) and ResNet50 (Right) trained for 100 epochs under a cosine learning-rate scheduler. Validation Top-1 errors (the inset highlights the late-epoch region) are reported.
As shown in Figure 5, Muon achieves faster convergence in the early phases, while KOALA++ initially exhibits larger fluctuations but gradually stabilizes and reaches comparable or better final accuracy. We attribute this difference to the distinct curvature modeling mechanisms of the two optimizers. Muon performs explicit matrix preconditioning with guaranteed positive semi-definite (PSD) curvature estimates, whereas KOALA++ relies on an implicit least-squares projection derived from scalar observations. Since the resulting covariance-like matrix is not explicitly constrained to be PSD, KOALA++ may require several epochs to implicitly correct unstable directions before converging to a well-conditioned update regime. This interpretation remains a hypothesis, and a rigorous theoretical analysis of the PSD property and its impact on convergence stability is left to future work.
Appendix D Ablations
D.1 Effect of Batch Size on KOALA++ Performance
We evaluated the impact of batch size on the performance of KOALA++ using ResNet50 on the CIFAR100 dataset. As shown in Table 6, increasing the batch size leads to a noticeable degradation in both Top-1 and Top-5 accuracy. The results suggest that KOALA++ is more effective in smaller-batch training regimes, possibly due to better gradient variance estimation and more frequent updates.
Table 6: Impact of batch size on KOALA++ performance (ResNet50, CIFAR100).
Batch Size Top-1 Error (%) Top-5 Error (%)
64 20.27 4.91
128 20.38 4.76
256 21.78 5.84
512 23.63 6.59
D.2 Different Schedulers
In this experiment, we evaluate the impact of different learning rate schedulers on the performance of KOALA++ using the CIFAR-100 dataset. We consider two representative architectures: ResNet-50 as a typical CNN-based model and Swin-Tiny as a typical ViT-based model. All models are trained for 200 epochs using the same optimizer (KOALA++) and hyperparameters.
The partially recorded results are summarized in Table 7.
Table 7: Top-1 and Top-5 error (%) on CIFAR-100 using KOALA++ with different schedulers.
Schedulers ResNet-50 (CNN) Swin-Tiny (ViT)
Top-1 Error Top-5 Error Top-1 Error Top-5 Error
Multi-step 21.80 5.89
CosineAnnealingLR 20.57 5.07 35.05 12.07
Warmup + Cosine 32.18 10.64
As Table 7 shows, different schedulers can lead to notable differences in final top-1 and top-5 error. We now take a closer look at how these schedulers influence the test loss dynamics during training.
For ResNet-50, we compare the Multi-step scheduler and the CosineAnnealingLR scheduler. As shown in Figure 6, CosineAnnealingLR results in a smoother and more stable descent in test loss, especially during the late training stages. This stable convergence leads to slightly better final performance compared to the traditional multi-step decay.
For the ViT-based Swin-Tiny model, we evaluate both CosineAnnealingLR and a Warmup + Cosine schedule. We find that adding a warm-up stage significantly enhances convergence speed and test loss. This is consistent with the behavior of transformer-based architectures, which are more sensitive to large initial learning rates. The Multi-step scheduler is not included in this setting, as it is generally less effective for transformers.
(a) ResNet-50: Multi-step vs. CosineAnnealingLR
(b) Swin-Tiny: CosineAnnealingLR vs. Warmup + Cosine
Figure 6: Log-scale test loss curves of KOALA++ on CIFAR-100 under different learning rate schedulers. Left: CosineAnnealingLR provides a smoother convergence than Multi-step on ResNet-50. Right: Warmup + Cosine improves stability and convergence for Swin-Tiny compared to CosineAnnealingLR alone.
Note: We observed that the Swin-Tiny model tends to overfit in the later stages when trained for 200 epochs. While we report the final Top-1 and Top-5 error at epoch 200 in Table 7 for consistency, the plotted curves in Figure 6 only show the first 100 epochs. This decision was made to focus on the informative convergence behavior and improve the visual clarity of the comparison.
D.3 Different Initializations of KOALA++
The initialization of the directional covariance product $v_{1}$ in KOALA++ depends directly on the initial covariance matrix $P_{0}$. To study the impact of $\sigma$ on the performance of the model, we conducted an ablation experiment in which we vary $\sigma\in\{0.05,0.1,0.2\}$ while keeping other hyperparameters fixed.
We perform this experiment on the CIFAR-100 dataset using the ResNet-50 architecture, training for 200 epochs. The learning rate is set to $2.0$, and the Kalman process noise parameter $q$ is fixed at $0.2$. The results are summarized in Table 8.
Table 8: Effect of varying $\sigma$ on KOALA++ performance. Dataset: CIFAR-100, Model: ResNet-50, Epochs: 200, $q=0.2$, lr=2.0=2.0
$\sigma$ Top-1 Error (%) Top-5 Error (%)
0.05 22.95 6.60
0.10 22.04 6.11
0.20 21.80 5.89
We observe that increasing $\sigma$ leads to improved performance, suggesting that initializing $v_{1}$ with stronger magnitude provides better early-stage directionality, which helps escape suboptimal regions in the optimization landscape. The results are shown in Figure 7.
(a) Log-scale Test Loss
(b) Top-1 Validation Error (%)
Figure 7: Effect of varying $\sigma$ on KOALA++ performance over 200 epochs on CIFAR-100. Larger $\sigma$ values yield more stable convergence and lower final error.
Appendix E Experiments Details
E.1 Hardware
All experiments reported in this work were conducted on a server equipped with a single NVIDIA H100 GPU with 80 gigabytes of VRAM and 128 gigabytes of RAM. Unless otherwise stated, all training and evaluation tasks were executed using this configuration.
E.2 Image Classification
E.2.1 HP Settings for CIFAR-10/100 and Hyperparameter Sensitivity
We evaluate KOALA++ on both CIFAR-10 and CIFAR-100 using a variety of architectures, including CNN-based models (e.g., ResNet [9]) and vision transformers (e.g., Swin-Tiny [18]). Across all experiments, we used the same hyperparameter settings for CNN-based models to highlight KOALA++’s low-tuning usability, as it achieves competitive or better results than baselines without dataset-specific tuning. When moving to Transformer-based architectures, we conducted a small grid search following the procedure described in the Supplementary Material.
Below we provide further clarification on how the key hyperparameters are selected, consistent with the analysis included in the Supplementary Material to support practical deployment of KOALA++:
  • Measurement noise $R$: estimated online as an exponential moving average of the mini-batch loss variance, using a smoothing factor $\alpha=0.9$. This removes $R$ from the set of hyperparameters requiring manual tuning.
  • Initial variance $\sigma_{0}$: controls only the scale of the initial directional state $v_{1}=\sigma_{0}H_{1}$. Ablation studies indicate that KOALA++’s convergence is robust to this value, so we fix $\sigma_{0}$ to match the process noise variance $Q$ for simplicity and symmetry.
  • Process noise $Q$: the main hyperparameter that requires tuning besides the learning rate. Empirically, we found that $Q\in[0.1,0.4]$ performs well across datasets.
For CIFAR-10, we initialize both $\sigma_{0}$ and $Q$ to $0.1$, with an initial learning rate of $1.0$. For CIFAR-100, which has more classes and a richer data distribution, we adopt slightly larger values $\sigma_{0}=Q=0.2$ and increase the initial learning rate to $2.0$. A weight decay of $5\times 10^{-4}$ is applied to all ResNet and other CNN models.
When applying KOALA++ to Swin-Tiny, we perform a coarse grid search over the weight decay parameter, evaluating values $\{1\times 10^{-1},1\times 10^{-2},5\times 10^{-3},1\times 10^{-4},5\times 10^{-4}\}$. Interestingly, KOALA++ performs best with a smaller weight decay ($1\times 10^{-4}$), whereas other optimizers typically favor higher decay values (e.g., $1\times 10^{-2}$). We hypothesize that this stems from KOALA++’s implicit regularization through dynamic covariance updates, which already stabilize parameter updates. Thus, combining KOALA++ with a large explicit regularizer (e.g., strong weight decay) can lead to underfitting—especially in transformer-based models—making a smaller weight decay more suitable in such cases.
E.2.2 More Results for CIFAR10/100
In the main body of the paper, we only visualized 100-epoch results on CIFAR-100 using ResNet-50 to compare KOALA++ with several baselines. Here, we provide more comprehensive results by extending the training to 200 epochs and including CIFAR-10 as well.
CIFAR10
Figure 8 (Left) presents the Top-1 validation error on CIFAR-10 with ResNet-50 across 200 training epochs. To highlight the behavior in the final stage, we also plot a zoomed-in view of the validation error during epochs 160–200 in Figure 8 (Right).
(a) Validation Error (Epoch 1–200)
(b) Validation Error (Epoch 160–200)
Figure 8: Validation Top-1 Error (%) on CIFAR-10 using ResNet-50. Left: Full 200-epoch curve. Right: Zoomed view of epochs 160–200.
Note:This figure shows one random seed; KOALA++ outperforms SGD on average across three runs.
CIFAR100
We additionally report the 200-epoch results on CIFAR-100 using ResNet-50, comparing KOALA++ with several standard optimizers. The log-scale test loss and Top-1 validation error are shown in Figure 9.
(a) Log-scale Test Loss
(b) Top-1 Validation Error (%)
Figure 9: Results on CIFAR-100 with ResNet-50 over 200 epochs. Left: Log-scale test loss. Right: Top-1 validation error.
E.2.3 Experiments for ImageNet32
We train both ResNet18 and ResNet50 models on the ImageNet32 dataset. Following standard augmentation practices for low-resolution image classification, we apply RandomCrop(32, padding=4) and RandomHorizontalFlip() during training. The input images are subsequently converted to tensors and normalized using the standard ImageNet statistics. For validation, we directly normalize the images without applying any cropping or resizing. All models are trained for 100 epochs with a learning rate scheduler that decays the learning rate by a factor of 0.1 at epochs 30, 60, and 90, and a weight decay of 0.0001.
Note: This setup differs slightly from the data augmentation strategy used in KOALA’s ImageNet experiments [4], where validation images are center-cropped prior to evaluation.
Figure 10 shows the log-scale test loss and Top-1 validation error for ResNet18. We observe that SGD reaches a lower test error faster in early training but tends to overfit slightly in the later stages. In contrast, KOALA++ shows improved generalization in the long run, despite being less stable early on.
Table 9 summarizes the final Top-1 and Top-5 errors for both ResNet18 and ResNet50. The results show that KOALA++ achieves comparable or better performance to SGD without requiring any task-specific hyperparameter tuning.
Table 9: Final Top-1 and Top-5 validation error (%) on ImageNet32.
Model Optimizer Top-1 Error Top-5 Error
ResNet18 SGD 45.81 22.02
KOALA++ 45.71 21.94
ResNet50 SGD 40.17 17.57
KOALA++ 39.03 17.12
Figure 10: Left: Log-scale test loss; Right: Top-1 validation error for ResNet18 on ImageNet32.
E.3 Efficiency Analysis
We provide additional results on the computational efficiency of KOALA++. Besides accuracy, it is important that an optimizer remains efficient in terms of runtime and memory, particularly when scaling to larger models and batch sizes.
ResNet-50 and Swin-Tiny (batch size 256).
Table 10 reports average training time per epoch, asymptotic FLOPs, and peak memory usage on ResNet-50 and Swin-Tiny. KOALA++ achieves runtime and memory comparable to Adam(W), while being substantially more efficient than Shampoo and AdaHessian. For K-FAC, we note that the more precise per-iteration complexity is $O(\ell d^{2}m)$ [^8] Here $\ell$ is the number of layers, $d$ the typical hidden dimension, and $m$ the batch size., although it is commonly reported as $O(n^{2})$ in the literature.
Table 10: Efficiency profiling on ResNet-50 and Swin-Tiny (batch size $256$). We report average training time per epoch, asymptotic FLOPs, and peak memory usage.
ResNet-50 (Batch size 256)
Optimizer Time/epoch (s) FLOPs (class) Peak Mem (GB)
Adam 11.08 $O(n)$ 10.51
AdaFisher 11.64 $O(n)$ 11.91
AdaHessian [^9] AdaHessian uses Hutchinson’s trick for Hessian diagonal estimation, which requires additional Hessian-vector products, leading to $O(n^{2})$ complexity. 63.43 $O(n^{2})$ 22.79
K-FAC [^10] More precisely $O(\ell d^{2}m)$, where $\ell$ is the number of layers, $d$ the hidden dimension, and $m$ the batch size. 17.16 $O(n^{2})$ 12.34
Shampoo 134.63 $O(n^{1.5})$ 10.41
KOALA++ (ours) 13.23 $O(n)$ 10.27
Swin-Tiny (Batch size 256)
Adam(W) 16.35 $O(n)$ 3.46
AdaFisher 12.00 $O(n)$ 4.38
AdaHessian 42.00 $O(n^{2})$ 5.97
KOALA++ (ours) 16.10 $O(n)$ 4.50
ResNet-101 and Swin-Base (batch size 1024).
To further test scalability, we report results on ResNet-101 and Swin Transformer Base with batch size $1024$ in Table 11. KOALA++ remains competitive with Adam(W), while scaling better than AdaHessian and Shampoo in both runtime and memory. This confirms that KOALA++ maintains near-first-order efficiency even at larger scales.
Table 11: Efficiency profiling on ResNet-101 and Swin Transformer Base with batch size $1024$.
ResNet-101 (Batch size 1024)
Optimizer Time/epoch (s) FLOPs (class) Peak Mem (GB) Adam 16.79 $O(n)$ 48.16 AdaFisher 16.62 $O(n)$ 48.11 K-FAC 23.79 $O(n^{2})$ 52.09 Shampoo 64.29 $O(n^{1.5})$ 48.90 KOALA++ (ours) 17.61 $O(n)$ 48.25
Swin Transformer Base (Batch size 1024)
Optimizer Time/epoch (s) FLOPs (class) Peak Mem (GB) Adam(W) 19.07 $O(n)$ 23.23 AdaFisherW 18.95 $O(n)$ 22.28 AdaHessian 57.72 $O(n^{2})$ 64.08 KOALA++ (ours) 19.52 $O(n)$ 22.51
Comparison with SGD across batch sizes.
We also compare throughput with SGD on CIFAR-10 using ResNet-50 under varying batch sizes (Table 12). KOALA++ incurs a small constant overhead per iteration due to $O(n)$ covariance–vector products, but this cost is independent of batch size. As the batch size grows, forward/backward cost dominates, and the relative overhead of KOALA++ shrinks, approaching parity with SGD at batch size $1024$.
Table 12: Throughput comparison of KOALA++ vs SGD across different batch sizes on CIFAR-10 with ResNet-50.
Batch Size SGD (s/epoch) KOALA++ (s/epoch) Ratio (KOALA++ / SGD) Overhead
[…]
This trend aligns with our design intuition: within each batch, the effective covariance dimension becomes smaller as stochastic noise is averaged out, so the extra operations of KOALA++ contribute less to the total runtime. Consequently, KOALA++ scales gracefully with batch size and approaches the efficiency of SGD in large-batch regimes.
E.4 Language Modeling
We adopt the experimental benchmark setup provided by AdaFisher [7], available at https://github.com/AtlasAnalyticsLab/AdaFisher. For a fair comparison, we use the same experimental pipeline and only test on the WikiText-2 dataset using the small GPT1 model.
To tune the hyperparameters of KOALA++, we performed a grid search over the following ranges:
  • Learning Rate (lr): {1e-1, 5e-2, 2e-2, 1e-2, 5e-3, 2e-3, 1e-3, 5e-4}
  • sigma and q (set equal): {0.01, 0.02, 0.05, 0.1}
  • Weight Decay: {1e-1, 5e-2, 2e-2, 1e-2, 5e-3, 2e-3, 1e-3, 5e-4, 2e-4, 1e-4}
The optimal configuration we found was: Learning Rate: 2e-3, sigma = q: 0.1, Weight Decay: 1e-4, which achieved the best perplexity among all tested settings. Figure 11 displays the validation loss and test perplexity curve of KOALA++ on the WikiText-2 dataset.
(a) Validation PPL vs Time (log)
(b) Validation Loss vs Time (log)
(c) Validation PPL vs Epoch (log)
Figure 11: Performance of KOALA++ on WikiText-2 with the small GPT1 model. Left: Log-scale validation perplexity over wall clock time. Right: Log-scale validation perplexity over training epochs. Following [7], AdamW was trained for 55 epochs, while KOALA++ were trained for 50 epochs.